### v3

# NonGEMM Bench: Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

Abstract
Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that construct the main backbone of ML models. As their computational overhead dominates the overall execution time (e.g., 49% - 94%in our results), GEMM-based operators have been the prime optimization target for fast ML inference and training, which led to advanced GPUs and accelerators available today. Such GEMM accelerators boost the GEMM performance and provide significant end-to-end speedup compared to CPUs, aligned with the lesson from Amdahl’s law. However, accelerating GEMM has shifted the amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant. Nonetheless, the performance of non-GEMM operators has not been studied with the latest GPUs. In addition, recent models include more complex non-GEMM operators (e.g. SiLU), which motivates the characterization of the performance to guide the future ML system design.
Therefore, we first extensively profile the execution time and energy of 18 widely adopted ML models in Hugging Face and Torchvision on mobile, workstation, and data center platforms with/without GPUs. Based on the results, we analyze the type of non-GEMM operators that contribute most to the execution time for each task. Finally, we collect all models with our performance profiling harness and create a new ML benchmark, NonGEMM Bench, which automatically reports detailed breakdown into GEMM and NonGEMM operators. NonGEMM Bench also provides a collection of 1460 non-GEMM operators with realistic input arguments and data to guide the non-GEMM operator-oriented system optimizations. Our case study shows the non-GEMM operators account for 60% of the total execution time, on average, across theree hardware platforms with GPUs. In addition, we report that the dominant non-GEMM operator varies with the task and even for different ML deployment software tools. Our results indicate that considering non-GEMM operator is important for future hardware and software system designs for ML systems.
1 Introduction
The success of machine learning (ML) in various problem domains, such as computer vision (CV)  [19, 18, 39, 25, 29] and natural language processing (NLP) [24, 5, 41], made ML workloads pervasive in various computing platforms from edge to cloud devices. ML model inference involves billions of multiply-and-accumulate (MAC) operations (e.g., 497 billions of MAC operations for ResNet 50 [19]). Such MAC operations originate from GEneral Matrix Multiplication (GEMM)-based operators, such as CONV2D, Linear, and BMM (batched matrix multiplication). The GEMM-based operators dominate in terms of the total execution time on CPUs, as shown in Figure 1. Therefore, GPUs and accelerators have focused on the optimization of the GEMM-based operators, which significantly enhanced the computational performance (e.g., latency and throughput) of end-to-end ML model inference, as highlighted in Figure 1.
Figure 1: The latency breakdown into GEMM and non-GEMM operators on AMD EPYC 7763 + NVIDIA A100 GPU. We measure the latency on two popular models from HuggingFace (a) GPT2-XL (batch 1) [5] and (b) Vision Transformer (batch 1) [25].
However, because the GEMM operators are being accelerated, the non-GEMM operators, such as memory operations (e.g., reshape, view, and transpose), normalization, and logit computation functions (e.g., Softmax), now account for a considerable amount of the end-to-end latency, compared to that of GEMM operators.  Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [5]) and image classification (Vision Transformer [25]) models. The motivational data imply a major shift of the landscape of Amdahl’s law [21] for ML acceleration, indicating that we now need to consider non-GEMM operators in the ML system optimization. However, no study thoroughly explored the non-GEMM operator performance in latest models, although it is important to quantitatively understand the current status due to the fast evolving GEMM accelerators such as GPUs.
Therefore, we propose NonGEMM Bench, which focuses on non-GEMM operator analysis in popular ML models today, and perform a thorough peformance characterization of Non-GEMM operators in the latest models widely adopted by the community. Unlike previous benchmarks, NonGEMM Bench thoroughly analyzes NonGEMM operator performance, providing important insights on the ML system optimization with powerful accelerators. NonGEMM Bench consists of 18 models collected on Hugging Face [43] and Torchvision [30] on various computer vision and natural language processing tasks.
To enable non-GEMM-oriented system optimization, NonGEMM Bench provides a microbenchmark of non-GEMM operators. In addition to the preset of non-GEMM operators collected in popular Hugging Face models, NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch FX  [37] and ONNX  [8], which provides flexibility to users. The flow allows users to analyze any non-GEMM operators without manually writing custom functions, unlike previous benchmark [27]. In addition, NonGEMM Bench utilizes input argument specification extracted from real data. This helps analyze non-GEMM operators in realistic settings.
Using this benchmark, we perform case studies on 18 popular Hugging Face and Torchvision models on image classification, object detection, segmentation, and language models (e.g., Llama 2 [41] and GPT-2 [35]). We study the performance on different hardware configuration ranging from mobile, workstation, to data center, which includes four CPU and four GPU models. In our case studies, we evaluate the effect of GPU acceleration on the relative latency across operators of NonGEMM Bench models and highlight that the contribution of non-GEMM operators to the end-to-end latency of the accelerated model.
We will open-source NonGEMM Bench at the time of publication to facilitate the ML system research aligned with the current performance horizon. Including that, we make the following contributions:
  • We shed a light on the changed landscape of Amdhal’s law in ML system design, which shows the increased importance of non-GEMM operators in systems with GEMM accelerations.
  • We propose NonGEMM Bench, which provides analysis on non-GEMM operators in the 18 latest popular models from Hugging Face and Torchvision.
  • To facilitate non-GEMM operator-aware optimizations, we organize a microbenchmark of non-GEMM operators collected in the popular models and provide realistic input arguments collected from realistic input data.
  • Using NonGEMM Bench, we perform case studies on three different hardware configurations ranging from mobile, workstation, and to data center, and show the non-GEMM operators are becoming major consideration across all platforms.
  • We identify different set of dominant non-GEMM operators depending on the task domain and deployment SW flow, which motivates domain-oriented specialization and shows the importance of deployment software’s role for non-GEMM operators.
2 Background
Figure 2: A description of Linear (a) and Conv1D (c) operators as GEMM operators example, and that of non-maximum suppression [18] (b) and Layer Normalization [7] (d) as a non-GEMM operator example.
To introduce NonGEMM Bench, we discuss the ML operators and recent ML models.
2.1 ML Operators
ML operators are the building blocks for ML models, which define the computation over one or multiple input tensors. Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization functions, as listed in Figure 2. Operators can be categorized into two classes: GEMM operation-based and the others, which we term as non-GEMM operators. We discuss each class of ML operators next.
2.1.1 GEMM-based Operators (GEMM Operators)
GEMM-based operators (or GEMM operators) refer to all the ML operators that can be represented as a matrix multiplication operation, which include linear, Conv2d, and batched-matrix multiplication (BMM). For example,  Figure 2 (a) and (c) illustrate two popular GEMM operators: Linear and Conv2d operators, respectively. Each operator can be represented into a perfectly nested loop with multiply-and-accumulate (MAC) operation in the inner-most loop. Note that variants that are not matrix multiplication in the default form like Conv2d can be converted into GEMM (e.g., im2col [13]), which motivated the term, GEMM operator.
GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in in Figure 1 show. However, they have regular computation patterns that can be summarized as a perfectly nested for loop. The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [33, 26, 46]
With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [14]) and accelerators [23].
2.1.2 Non-GEMM Operators
Non-GEMM operators refer to all ML operators other than GEMM operators. They span various functionalities (e.g., memory layout manipulation and normalization) other than applying weights to input tensors. Because of their diverse functionalities, their computation patterns are often not a perfectly nested loop with MAC, which can also involve non-linear functions and memory-oriented operations. For example,  Figure 2 (b) shows non-maximum suppression (NMS) operator often found in R-CNN model variants [18, 38]. As found in the example, the entire operation cannot be summarized into single perfectly-nested loop, which involve other operations such as sort and filtering. In addition, the operation involves a conditional statement, which introduces non-deterministic behaviors to the operator. The layer normalization example in Figure 2 (b) also shows another key characteristic of the non-GEMM operators: non-linear functions. Because of such characteristics distinguished from GEMM operators, optimization methodologies for GEMM operators cannot be applied to accelerate non-GEMM operators.
To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 18 recent models in the computer vision and natural language processing domains. We select models based on their popularity in the Hugging Face community to obtain realistic workload. We list the models we investigated in Table 1. Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table 2.
  • Normalization. Normalization operators regularize the data range across a selected dimension using the mean and standard deviation statistics. Examples include BatchNorm [22] and LayerNorm [7], which are widely adopted in computer vision and NLP ML models [18, 5]. Such variants defer by the choice of the normalized dimension (e.g., BatchNorm normailizes a tensor across the batch dimension). RMS Norm [47], which is adopted in recent large language models [41], is another example of the normalization function. RMS Norm eliminates the division by standard deviation in typical normalization functions and performs $\sqrt{\frac{1}{n}\Sigma^{n}_{i=1}(X_{i}-\mu)}$, where $X_{i}$, n, and $\mu$ refer to the i-th data, number of data, and the mean, respectively.
  • Activation. Activation operators introduce non-linearity into the model. Rectified Linear Unit (ReLU) function [32] is an example of activation operators used in CNN based ML models  [39, 19, 40]. It injects non-linearity into the model based on the sign of the data by applying the function $X_{i}=Max(0,X_{i})$ element-wise to the data, where $X_{i}$ is the i-th data point. Another variant of activation operators is the the Gaussian Error Linear Units function (GELU) [20] which is a popular activation function adopted in transformer based ML models s [5, 45, 25, 29]. Unlike ReLU, GELU accounts for the value of the data when inserting the non-linearity and not only the sign. It multiplies the input $X_{i}$ by the Cumulative Distribution Function (CDF) of a Gaussian distribution: $GELU(X_{i})=X_{i}*\phi(X_{i})$ [20] where $X_{i}$ is the i-th data and $\phi$ is the CDF of the Gaussian distribution.
  • Memory Operators. Memory operators are responsible for the allocation and the layout modification of tensors inside the memory. Examples cover the reshape and concatenation (cat) operators. Reshape operators do not create new tensors. They modify the shape (e.g. dimension order) of a tensor and return a new view of the tensor following the new dimension order. While reshape operators do not allocate additional memory, the cat operator allocates new memory to store a newly created tensors by concatenating the input tensors along a new dimension.
  • Arithmetic. Arithmetic operators cover non-GEMM arithmetic operators applied on tensors in an element-wise or reduction pattern. Element-wise division applied to scale the elements of tensors in the attention block as shown in Figure 3 (c) is an example of such operators.
  • RoI Selection. RoI selection operators are found in R-CNN variant detection models. [38, 18]. They filter down bounding boxes proposed by the region proposal network (Figure 3 (b)) and align the remaining boxes to the objects detected in the image. Non-Maximum Suppression (NMS) is an example of RoI Selection, which is described in Figure 2. Given a list of scores and bounding box information, it selects bounding boxes by applying the Intersection over Union (IoU) metric.
Figure 3: Architectures of three popular ML model families.
2.2 ML Models and Popular Tasks
The heterogeneity in non-GEMM operators enabled ML engineers and researchers to build models supporting a wide range of tasks (e.g. Computer Vision and NLP). As highlighted in  Figure 3, computer vision ( Figure 3 a and b) and NLP ( Figure 3) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
For example, classical models in image classification shown in  Figure 3 (a) consist of deep convolution networks (CNN). which cascade Conv2d (GEMM) with normalization and activation operators [39, 19]. Object Detection models also rely on convolution network for feature extraction, region proposal and classification. However, they pair the CNNs with ROI selection non-GEMM operators to process and filter the detections (Figure 3 (b)). On the other hand, NLP and language models rely on the transformer architecture, which leverages the attention mechanism introduced in [42]. Transformers combine BMMs and Linear GEMM operators with normalization, memory and arithmetic non-GEMM operators as shown in Figure 3. This diversity in non-GEMM operators and model architectures across tasks paired the preliminary data presented in  Figure 1 motivate us to study the impact of non-GEMM operators have on the inference performance of popular models.
Figure 4: An overview of NonGEMM Bench flows
3 NonGEMM Bench
Understanding the performance landscape of ML models with non-GEMM workloads necessitates a specialized benchmark. A good benchmark targeting non-GEMM operators should 1) capture the operator level performance breakdown not only in a standalone scenario, but also in an end to end inference to understand the effects of the non-GEMM operators. Moreover, it should analyze 2) frequently used models in 3) diverse task domains on 4) real data to reflect the actual performance of realistic application.
ML Benchmarks available today (e.g., MLPerf [36]), unfortunately, do not satisfy the requirements since they do not focus on the non-GEMM operators. Long-tail bench [27] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domain. To better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench, that satisfies all the requirements for a non-GEMM operator specific benchmark. NonGEMM Bench provides operator-level breakdown of execution time at the computational/operator graph level in end-to-end inference scenarios, which enables detailed non-GEMM operator performance analysis (e.g., impact of each operator type on the end-to-end performance), as we present in Section 4. In addition, to facilitate non-GEMM operator-focused system optimizations, NonGEMM Bench includes a micro-benchmark of non-GEMM operators with concrete input arguments and tensors collected from realistic datasets and models. We discuss the datasets and models adopted in NonGEMM Bench and describe the structure of NonGEMM Bench in detail next.
3.1 NonGEMM Bench Models
To capture the performance in the latest ML workload, we select 18 highly downloaded models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench.  Table 1 lists the NonGEMM Bench model registry (Figure 4) which contains 18 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench. The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
3.1.1 Computer Vision (CV)
Deep Learning is widely used in the CVdomain [25]  [19]  [38]. NonGEMM Bench covers three important tasks of the domain:
Image Classification (IC). Image classification refers to a CV task that identifies a class label of a given image. Image classification models extract features (i.e., high-level and -dimensional information of the input image) from an input image and reports the class label utilizing the features. NonGEMM Bench includes eight most popular IC models in HuggingFace [1], which includes ResNet50  [19], MobileNetv2  [39], three variants of Vision Transformer [25], and three variants of Swin Transformer  [29].
Image Segmentation (IS). Image segmentation refers to a computer vision task that identifies the area in a image for each class. Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels belong to each object. NonGEMM Bench includes two state-of-the-art IS models: Segformer  [45] and MaskFormer  [12]. Object Detection (OD). Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object. Like aforementioned models, OD models extract features. In addition to that, OD models generate region proposal, which refers to the candidate location and bounding boxes for objects in an image. Using Region of Interest (RoI) processing non-GEMM operators listed in Table 2, OD models refine the raw region proposal generated by a region proposal network. The refined RoIs are used as inputs to the classifier CNN at the end, and the classifier determines the class label of object inside each refined RoI. NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [38], MaskRCNN  [18], and DETR  [10].
3.1.2 Natural Language Processing (NLP)
NLP refers to tasks involving the analysis and understanding of human (natural) language. NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [48] [9]. Transformer [42] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state of the art generative large language models like GPT  [35] and Llama  [41].  Figure 3 (c) shows the layer architecture of GPT’s transformer. It consists of a self attention block build by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table 2). NonGEMM Bench includes five popular NLP models  [3]: Bert [24], three variants of GPT2 [5], and Llama2-7b [41].
3.2 Benchmark Suite
NonGEMM Bench provides two flows: end-to-end inference and microbench flows. The End-to-end flow evaluates the performance of operators when the model is running inference on dataset data. The Microbench flow performs micro-benchmarking on the operator by running it in a standalone fashion using synthetic data with realistic dimensions. We discuss input, output, and the benchmark software structure for each flow.
3.2.1 NonGEMM Bench Inputs
NonGEMM Bench receives target deployment software flow and other configurations as inputs. Optionally, users can provide custom models with associated datasets to profile their workloads. By default, NonGEMM Bench runs the 18 models in Table 1. We describe each input in detail as follows:
Models. As described in  Section 3.1, NonGEMM Bench offers a registry of pre-selected popular ML models for users to analyze. Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators. Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
Deployment Flow. NonGEMM Bench deploys its models using two popular inference frameworks: ONNX Runtime [16] and PyTorch [34] eager mode, TorchScript, and TorchDynamo introduced in PyTorch 2.0 [6].
Datasets. To evaluate the models, NonGEMM Bench utilizes real data from popular datasets in each domain. We use ImageNet 2012  [15] and MS COCO  [28] for computer vision tasks. As for language models, we use wikitext dataset [31] available on HuggingFace [4].
Misc. Configurations. This includes all other configurations for profiling runs, which includes the batch size, the number of profiling iterations, and the target device.
Table 1: Tasks and Models in NonGEMM Bench
Application Models # Params Dataset
Image Classification ResNet50 [19] 0.85M ImageNet [15]
MobileNetv2 [39] 3.4M
ViT large 16 [25] 307M
ViT huge 14 [25] 632M
Swin tiny [29] 29M
Swin small [29] 50M
Swin base [29] 88M
Object Detection FasterRCNN [38] 42M COCO [28]
MaskRCNN [18] 44M
DETR [10] 41M
Segmentation Maskformer [12] 102M COCO [28]
SegFormer [45] 3.7M
Language Models GPT2 [5] 117M wikitext [31]
GPT2 Large [5] 762M
GPT2 X-Large [5] 1.5B
Llama 2-7B [41] 7B
Bert [24] 110M
3.2.2 End-to-End Inference Flow
The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset, which consists of two stages. The first stage extracts the operator graph of a given model and preprocess input data. The second stage records the latency and energy of the inference runs.
Graph Extractor functions convert the model to be evaluated into a computational graph exposing the GEMM and non-GEMM operators for profiling. Based on the selected deployment flow, the graph extractor calls the corresponding APIs to trace through the model and generate the operator graph. For the ONNX Runtime deployment flow, the NonGEMM Bench harness calls ONNX graph exporters from the HuggingFace transformers library [44] or from the PyTorch libraries. For the PyTorch deployment flow, the graph extractor relies on the torch.fx [37] Tracer to obtain the computational graph. The Tracer propagates a proxy data structure into the model and captures the operators as graph nodes.
Data Preprocessing functions are model specific functions that feed the profiling functions with input tensors. First, they fetch the raw data from the specified dataset and perform data cleaning. Empty sequences are removed from text data and images are re-scaled to each model’s specific input resolution. Second, they apply tokenization or image to tensor transformation to generate the input tensors.
End-to-End Inference Profiling Functions receive operator graphs and input tensors from the harness functions and run profiling iterations. The latency profiling functions record the operator breakdown of the end to end latency as well as the shape of the operators input tensors. Similar to the graph extractors, the latency profiling functions are deployment flow specific. For ONNX Runtime, we invoke Execution Provider profiling in the inference session. Currently, we only support operator level profiling for CPU and CUDA execution providers because profiling support to other ONNX RUNTIME backends is not stable yet. In PyTorch eager mode, we modify the torch.fx Interpreter to invoke the PyTorch profiler at each node of the operator graph. In TorchScript and TorchDynamo, we currently rely on the PyTorch Profiler. The energy profiling functions record the end to end inference power and energy consumption. We utilize system specific profiling tools to obtain CPU and GPU power profiles: NVIDIA’s System Management Interface (nvidia-smi) and AMD micro-Prof ($\mu$Prof) software profiling analysis tool.
3.2.3 MicroBench Flow
The MicroBench flow focuses on individual non-GEMM operators, which allows users perform deep analysis of their systems targeting non-GEMM operators. The flow consists of MicroBench Extractor, the NonGEMM Bench Operators Registry, and the MicroBench Operator Profiling (shown as ”u-Bench op Profiling in Figure 4).
MicroBench Extractor. This component populates the non-GEMM operators captured during end-to-end flow of NonGEMM Bench to the microbench operator registry. In addition to the operators, the registry stores the shape of input tensors to the non-GEMM operators as well as the implementation in their parent model. Such a step is crucial for keeping the microbenchmark realistic.
MicroBench Operator Profiling function (”u-Bench op Profiling” in Figure 4) fetch the non-GEMM operators from the registry and profile their performance utilizing latency/energy profiling in the end-to-end flow. For each operator, it generates input tensors using the recorded shapes and proceeds to evaluate the latency and the energy of the operator.
3.2.4 NonGEMM Bench Outputs
NonGEMM Bench generates many statistics orgainzed into three categories: performance/cost, workload, and non-GEMM-specific reports. The performance/cost report includes key performance metrics such as the end-to-end latency with operator level break-downs, energy, and peak memory usage. The workload report includes the shape of the tensors captured during inference on realistic data. Finally, the non-GEMM report provides insights on non-GEMM operators such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework), non-GEMM operator trace on different domains, and so on.
Table 2: Non-GEMM operators in eight popular model variants selected from Hugging Face listed in Table 1.
Operator Group Operator Model Single Operation Single Operand Non Linearity Dynamicity Reduction Example Input Shape
Activation ReLu DETR [2,64,533]
GELU ViT-l16 [1, 97, 4096]
GELU GPT2-XL [1, 8, 6400]
SiLu Llama-2 [1, 10, 11008]
Normalization LayerNorm Segformer [2, 16384, 32]
BatchNorm2d Segformer [2, 256, 128, 128]
LlamaRMSNorm Llama [1, 10, 4096]
FrozenBatchNorm2d MaskRCNN [1, 1024, 50,68]
FrozenBatchNorm2d DETR [2, 850, 256]
LayerNorm DETR [2, 850, 256]
Arithmetic Add Segformer [2, 16384, 32]
Mul Llama-2 [1, 10, 11008]
Neg Llama-2 [1, 32, 10, 64]
TrueDiv Segformer [2, 1, 16384, 256]
TrueDiv GPT2-XL [1, 25, 8, 8]
Memory Contiguous Segformer [2, 32, 128, 128]
Contiguous Llama-2 [1, 10, 32, 128]
Permute ViT-b16 [1, 768, 196]
Permute GPT2-XL [1, 8, 25, 64]
Split GPT2-XL [1, 8, 4800]
View GPT2-XL [1, 8, 1600]
Reshape ViT-b16 [1, 768, 14, 14]
Expand ViT-b16 [1, 1, 768]
Squeeze Llama-2 [1, 1, 10, 128]
Logit Softmax DETR [1, 25, 8, 8]
Computation Softmax Segformer [2, 1, 16384, 256]
RoI Selection NMS MaskRCNN [4663, 4]
Interpolation Interpolate Segformer [2, 256, 128, 128]
4 Case Studies
We perform a thorough analysis of the performance data generated by running NonGEMM Bench on multiple hardware platforms show in  Table 3. We collect the data by running NonGEMM Bench in PyTorch eager mode on three systems configuration ranging from mobile CPUs to data center grade CPUs and GPUs to capture the behavior of the operators when subjected to various acceleration capabilities. We notice that for most of the models, the time spent to compute non-GEMM operators increases noticeably with respect to the total execution time when running the model on the GPU.
Table 3: Hardware platform configurations used for NonGEMM Bench case studies.
Class CPU GPU
Data Center AMD EPYC 7763 NVIDIA A100
Workstation Intel i9-13900K NVIDIA RTX 4090
Mobile Intel i7-13700HU NVIDIA RTX 4060m
4.1 Impact of Non-GEMM on Each Task Domain
We discuss how non-GEMM operator’s contribution to the end-to-end inference and energy changes depending on task domains.
4.1.1 Image Classification
As shown in Figure 5, we observe a clear change in relative break-down latency across operators when using a GPU. For example, the normalization operation is minor before GPU acceleration, but it becomes the dominant operator with a GPU, accounts for over 50% of the total execution time, on average. We observe similar patterns for all the other cases: GEMM operators is dominant on CPU-only systems, but non-GEMM operators become significant or considerable after GPU acceleration. For the batch size of 1, the relative execution time of Non-GEMM operators for running ViT-b16 and ViT-l16 increases from 23.5% and 18% when running on a CPU only platform to 60% and 55%, respectively, on the GPU as shown in  Figure 5. Similarly, for the workstation GPU system, the non-GEMM operators execution time proportion increases to 56% and 45% for ViT-b16 and ViT-l16 to almost split the total execution time evenly with GEMM operators. For the three Swin transformers with a batch size of 1, the non-GEMM operators contribution to the overall latency increases as well when running the models on the GPU. The small Swin transformer model spends 55% of the computation time executing non-GEMM operators on the data center GPU system compared to a modest 33% contribution on the CPU only system. On the workstation GPU system, the small Swin transformer also computes non-GEMM operators for 57% of the time, a 2x increase compared to the workstation CPU system.
Impact of the batch size. When increasing the batch size, the latency distribution across GEMM and non-GEMM operators of IC models running on GPU follows the same trend as that of single batch. However, this trend is less noticeable for the largest vision transformer, ViT huge (632M parameters). For a batch size of 8, the non-GEMM operator latency is 43% and 51% for ViT-l16 and base Swin transformer respectively (Figure 5) running on the data center GPU, an 3.9$\times$ and 3$\times$ increase compared to the CPU only results. However, the huge vision transformer only shows a 1.25x increase in the latency proportion of non-GEMM operators when accelerating with a GPU.
The non-GEMM operators with the most significance to the execution time of classification models are presented in Figure 5. As shown in  Table 4, Normalization operators are the most expensive non-GEMM operator in classification models accounting for an average of 17% of the execution time on the data center GPU, followed by Arithmetic operators with 14% (Figure 5). Based on NonGEMM Bench collected data, image classification models require  48% of execution time on non-GEMM operators on GPU systems, on average, across all the evaluated models.
4.1.2 Object Detection
Non-GEMM operator’s latency is also noticeable for object detection with a GPU. On the data center, non-GEMM operators dominated the execution time of FasterRCNN by consuming 82% of the total inference time when we accelerate the workload compared to 46% when running the CPU only (Figure 5). DETR and MaskRCNN showed a similar behavior after acceleration, with a 2x increase in the time spend on non-GEMM operators. DETR computed non-GEMM operators for 77% of the inference time on the data center GPU, while MaskRCNN spent 80% on the data center and workstation GPUs. For object detection models, Normalization operators exhibit the highest execution time followed by activation and memory operators as indicated in (Figure 5). All detection models in NonGEMM Bench use normalization functions with custom implementations that do not leverage available library normalization kernels. This leads to additional overhead to execute the normalization function. Overall, the evaluation data shows that detection models get dominated by non-GEMM operators (78%) when executing the workloads on a GPU system.
4.1.3 Segmentation
For segmentation models, the non-GEMM operators account for 47% of the execution time when running the inference on the CPU on average. Upon running the model on the GPU, this number slightly increases to 53%. Figure 5 shows memory and arithmetic operators to be the most latency expensive non-GEMM operator on GPU accounting for 17% and 12.5% respectively. For the SegFormer model, Normalization operators are the most expensive (15% of the execution time), while Memory operators dominate the non-GEMM operators of Maskformer (24%).
4.1.4 Language Models
GPT2 and Bert executed non-GEMM operators for 77% and 57% of the time when running on a single batch as indicated in Figure 5. For GPT2-large, the non-GEMM operators clearly dominated the execution time on GPU systems as well. Across the three GPU HW classes, non-GEMM operators occupy 76% of execution time on average. Table 4 show that Activation (23% on GPT2) and element wise arithmetic (20% on Bert) operators are the most expensive non-GEMM operators in language models running on GPU.
Table 4: NonGEMM Bench Most Expensive NonGEMM Operator Group for selected NonGEMM Bench Models Running on Data Center GPU Platform.
Models Model Alias Batch Size Operator Group % of Exec. Time
ViT Base 16 Vt-b 1 Norm 20
8 Norm 17
ViT Large 16 Vt-l 1 Norm 20.5
8 Arith 17
Swin Tiny Sw-t 1 Norm 19
8 Norm 18.5
Swin Small Sw-s 1 Norm 18
8 Norm 17
Swin Base Sw-b 1 Norm 18
8 Norm 18
FasterRCNN FRCNN 1 Norm 60.5
2 Norm 59
8 Norm 49
MaskRCNN MRCNN 1 Norm 58
2 Norm 56
8 Norm 49
DETR DETR 2 Norm 40
GPT2 gpt2 1 Act 23
64 Act 16
GPT2 X-Large GPT2-XL 1 Act 23
64 Act 9
Llama2-7B Llama2 1 Arith 23
Bert bert 1 Arith 20
64 Arith 10
Figure 5: Breaking down the execution time of models running on the Data Center (CPU only) and (CPU + GPU) configurations across operator groups.
Figure 6: Breaking down the execution time of models running on the Workstation (CPU only) and (CPU + GPU) configurations across operator groups.
Large language model’s latency is dominated by GEMMs when running on Data center, consisting of 79% of the total inference time on average across platforms. When utilizing the GPU, the GEMM latency no longer bottlenecks the total execution time. The GEMM operators time drops to 30% and 22% for Llama2-7b and GPT2-XLarge respectively with non-GEMM operators executing for 74% on the GPU configuration. Figure 7 (a) suggests that the bottleneck are activation, arithmetic, and memory operators. Element-wise arithmetic operators are the most time-consuming non-GEMM operators for Llama2-7b, accounting for 23% of the total runtime (Table 4, while activation functions occupy 23% of GPT2-XL total execution time.
GPT2 models implement their own GELU activation in the Hugging Face library which does not have a direct mapping to a GELU GPU kernel when running in Pytorch eager mode. This leads to additional overheads due to running multiple kernels to compute the activation. Similarly, Llama uses a normalization function that does not map to a GPU kernel in eager PyTorch which leads to similar overheads that increase the latency of the non-GEMM operators.
4.2 Impact of Deployment Flow on Non-GEMM
Running in eager PyTorch, Non-GEMM operators occupy large proportions of the execution time for various applications running on the GPUs , as summarized in Table 4. Nevertheless, deployment frameworks (e.g. [16, 11]) are typically used to improve the performance beyond the eager mode [6]. They apply transforms to the computational graph level and leverage accelerators (if available) to optimize the inference runs. To understand how deployment frameworks affect inference performance distribution across operators, we evaluate NonGEMM Bench models with ONNX Runtime (ORT). ORT is a cross platform ML framework that accelerates ML model inferencing [16]. It leverages hardware accelerators (e.g. GPUs) and operator graph level optimizations to improve the performance of ML models. We run selected models with the CUDA execution provider and observe the following results.:
Non-GEMM operators account for a significantly larger proportion of the total execution latency (Figure 8) when running on different grade GPUs. This is aligned with the trend observed with the PyTorch eager execution on GPU. On average, the time spend on non-GEMM operators increases from 52% to a 73% after deploying the models on GPU with the ORT cuda execution provider.
Memory operators almost dominate the execution time in three application domains, classification, detection and language models by occupying 66%, 44%, 59 % respectively. For instance, memory operators are the most latency expensive non-GEMM operators in GPT2-XLarge and Llama2-7b as they consume 76% and 59% of the total inference time respectively (Figure 7). In detection models, memory operators consume 37% of the total inference time in ORT as well (Figure 8).
One reason leading memory operators to dominate the inference latency is their high frequency, notably in the large language models. They consist of 80% and 62% of the total operators in the GPT2-XLarge and Llama2-7b . Consequently, the cumulative latency to run these operators becomes the inference latency bottleneck. Another reason is the limitations in operator support for GPU deployment in ORT. Memory operators are a subset of non-GEMM operators that the cuda execution provider does not fully support, hence it assigns them to be computed on the CPU. In the case we hit an unsupported operator while executing on the GPU, the runtime falls back to CPU to carry the computation. As a result, additional overheads to transfer the data between devices are added to the operators execution time that increases the latency of the fallen-back operators, which aggravates frequently used operators.
Despite dominating the execution time in ORT, memory operators were not the main non-GEMM bottlenecks PyTorch eager mode for the evaluated models. Figure 7 highlights the differences in operator latency distribution for GPT2-XLarge and Llama2-7b running in PyTorch eager mode and ORT. This indicates that the deployment flows have distinct effects on the execution of non-GEMM operators in an end to end inference.
Figure 7: The impact of deployment software toolchain into the latency breakdown. (a) PyTorch (Eager Mode) and (2) ONNX Runtime on a data center class GPU (A100).
Figure 8: The latency breakdown of the inference using ORT on two platform configurations: mobile and Data Center.
4.3 Key Observations and Insights
We summarize our main observations and insights:
  • Accelerating GEMM operators using GPUs significantly shifts the performance horizon, which makes the non-GEMM operators main bottleneck of the systems: On average, non-GEMM breakdown increases from 27% (CPU-only) to 55% (CPU + GPU).
  • We observe the dominance of non-GEMM operators after GPU acceleration across all the platforms (mobile, workstation, and datacenter) and task domains.
  • The choice of deployment software flow affects the most time-consuming non-GEMM operators: In ORT, memory operators dominate the execution time across all models (on average, 56%), while PyTorch doesn’t show clear patterns.
  • Dominant non-GEMM operator class vary depending on the task category: In IC models, normalization operators dominate (on average, 18.4% of the total execution time), while activation and element-wise arithmetic operators dominate in language models (17.75% and 17.6% for each, on average).
  • The profiling results indicate that the non-GEMM operators are now the prime optimization target after GEMM acceleration.
5 Related Works
Table 5: Comparing NonGEMM Bench with existing benchmarks: MLPerf [36], LongTail Bench [27], TorchBench [17], and NonGEMM Bench (This work).
Benchamrk MLPerf LongTailBench TorchBench NonGEMMBench (This work)
Real Usage Driven
Non-GEMM Focused
Real Dataset Driven
Plug Model & Profile
Based on the characteristics of non-GEMM operators, and on the results shown in  Section 4, we clarify different features of key related works comparing against this work in Table 5. TorchBench [17] presents a comprehensive benchmark suite for the Pytorch framework consisting of 84 models whose selection was driven by real usage like NonGEMM Bench. MLPerf [36] is an industry standard state-of-the-art inference benchmark, which focuses on the end-to-end performance. LongTail Bench [27] is a micro-benchmark of NonGEMM operators. However, LongTail bench is a microbenchmark that only focuses on custom implementations of 34 operations, which cannot accept new operators nor profile various implementation of the same operator. As summarized in Table 5, all work shed a light to individual important aspect, but only NonGEMM Bench provides all the desired features for a non-GEMM operator-focused benchmark.
6 Conclusion
Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law in favor of non-GEMM operators. To guide future optimization, we need to understand the performance of non-GEMM operators in recent popular models. In this work, we analyze the performance non-GEMM operators in recent popular machine learning models. We construct NonGEMM Bench to study of non-GEMM operators in accelerated popular models reflecting realistic use scenarios and evaluate it on different hardware platforms. We show that the dominance of non-GEMM operators is occurring in all platforms with GPUs, and the deployment software’s impact is large. Our findings strongly motivate ML system researchers to consider non-GEMM operators as a major consideration.

### v4

@@
-NonGEMM Bench: Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads
+Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads
 Abstract
-Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that construct the main backbone of ML models.
-As their computational overhead dominates the overall execution time (e.g., 49% - 94%in our results), GEMM-based operators have been the prime optimization target for fast ML inference and training, which led to advanced GPUs and accelerators available today.
-Such GEMM accelerators boost the GEMM performance and provide significant end-to-end speedup compared to CPUs, aligned with the lesson from Amdahl’s law.
-However, accelerating GEMM has shifted the amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant.
-Nonetheless, the performance of non-GEMM operators has not been studied with the latest GPUs.
-In addition, recent models include more complex non-GEMM operators (e.g.
-SiLU), which motivates the characterization of the performance to guide the future ML system design.
-Therefore, we first extensively profile the execution time and energy of 18 widely adopted ML models in Hugging Face and Torchvision on mobile, workstation, and data center platforms with/without GPUs.
-Based on the results, we analyze the type of non-GEMM operators that contribute most to the execution time for each task.
-Finally, we collect all models with our performance profiling harness and create a new ML benchmark, NonGEMM Bench, which automatically reports detailed breakdown into GEMM and NonGEMM operators.
-NonGEMM Bench also provides a collection of 1460 non-GEMM operators with realistic input arguments and data to guide the non-GEMM operator-oriented system optimizations.
-Our case study shows the non-GEMM operators account for 60% of the total execution time, on average, across theree hardware platforms with GPUs.
-In addition, we report that the dominant non-GEMM operator varies with the task and even for different ML deployment software tools.
-Our results indicate that considering non-GEMM operator is important for future hardware and software system designs for ML systems.
-1 Introduction
+Among ML operators today, GEneralMatrix Multiplication (GEMM)-based operators are known to be key operators that build the main backbone of ML models.
+As their computational overhead dominates the overall execution time (e.g., 42.8% - 96.6% in our results), GEMM operators have been the prime optimization targets for fast ML inference.
+This led to advanced GPUs and accelerators available today, which provided significant boost in the GEMM performance compared to CPUs, aligned with the lesson from Amdahl’s law.
+However, accelerating GEMM has significantly shifted the Amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant.
+Although the importance of non-GEMM performance is increasing, we have little knowledge about the non-GEMM performance horizon in the latest hardware platforms and models.
+Therefore, to guide non-GEMM-oriented optimizations, we conduct a thorough performance analysis of 16 widely adopted ML models in Hugging Face and Torchvision on workstation and data center platforms with/without GPUs.
+We discover that non-GEMM performance bottleneck is a considerable issue across all the platforms and models, accounting for 11.3% to 73.6% of total latency, on average.
+The challenge significantly aggravates when we apply quantization, which is a common model compression technique, due to the boosted GEMM performance and extra non-GEMM operators for dequantization and requantization.
+To provide insights into non-GEMM optimization targets, we demystify the most dominant non-GEMM operators for each model and deployment software.
+We also show that widely adopted optimizations such as operator fusion do not completely address the non-GEMM performance bottleneck, where non-GEMM operators still account for 15% to 48% of total latency.
+We will open-source our non-GEMM-oriented benchmark framework to facilitate research in non-GEMM optimization.
+I Introduction
 The success of machine learning (ML) in various problem domains, such as computer vision (CV)  [22, 21, 54, 32, 37] and natural language processing (NLP) [31, 4, 58], made ML workloads pervasive in various computing platforms from edge to cloud devices.
@@
 Figure 1: The latency breakdown into GEMM and non-GEMM operators on AMD EPYC 7763 + NVIDIA A100 GPU.
-We measure the latency on two popular models from HuggingFace (a) GPT2-XL (batch 1) [5] and (b) Vision Transformer (batch 1) [25].
+We measure the latency on two popular models from HuggingFace (a) GPT2-XL (batch 1) [4] and (b) Swin Transformer (batch 1) [37].
 However, because the GEMM operators are being accelerated, the non-GEMM operators, such as memory operations (e.g., reshape, view, and transpose), normalization, and logit computation functions (e.g., Softmax), now account for a considerable amount of the end-to-end latency, compared to that of GEMM operators.
-Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [5]) and image classification (Vision Transformer [25]) models.
-The motivational data imply a major shift of the landscape of Amdahl’s law [21] for ML acceleration, indicating that we now need to consider non-GEMM operators in the ML system optimization.
-However, no study thoroughly explored the non-GEMM operator performance in latest models, although it is important to quantitatively understand the current status due to the fast evolving GEMM accelerators such as GPUs.
-Therefore, we propose NonGEMM Bench, which focuses on non-GEMM operator analysis in popular ML models today, and perform a thorough peformance characterization of Non-GEMM operators in the latest models widely adopted by the community.
-Unlike previous benchmarks, NonGEMM Bench thoroughly analyzes NonGEMM operator performance, providing important insights on the ML system optimization with powerful accelerators.
-NonGEMM Bench consists of 18 models collected on Hugging Face [43] and Torchvision [30] on various computer vision and natural language processing tasks.
-To enable non-GEMM-oriented system optimization, NonGEMM Bench provides a microbenchmark of non-GEMM operators.
-In addition to the preset of non-GEMM operators collected in popular Hugging Face models, NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch FX  [37] and ONNX  [8], which provides flexibility to users.
-The flow allows users to analyze any non-GEMM operators without manually writing custom functions, unlike previous benchmark [27].
-In addition, NonGEMM Bench utilizes input argument specification extracted from real data.
-This helps analyze non-GEMM operators in realistic settings.
-Using this benchmark, we perform case studies on 18 popular Hugging Face and Torchvision models on image classification, object detection, segmentation, and language models (e.g., Llama 2 [41] and GPT-2 [35]).
-We study the performance on different hardware configuration ranging from mobile, workstation, to data center, which includes four CPU and four GPU models.
-In our case studies, we evaluate the effect of GPU acceleration on the relative latency across operators of NonGEMM Bench models and highlight that the contribution of non-GEMM operators to the end-to-end latency of the accelerated model.
-We will open-source NonGEMM Bench at the time of publication to facilitate the ML system research aligned with the current performance horizon. Including that, we make the following contributions:
+Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [4]) and image classification (Swin Transformer [37]) models.
+The motivational data imply a major shift of the landscape of Amdahl’s law [24] for ML inference acceleration, indicating that we now need to consider non-GEMM operators as one of the major optimization targets in the ML system optimization.
+However, the research community today lacks a thorough and systematic performance analysis and characterization of non-GEMM operators in the latest models, which hinders the development of non-GEMM oriented optimization techniques.
+Therefore, we collect widely-used ML models from Hugging Face [61] and Torchvision [38] and perform a thorough performance characterization of of non-GEMM operators in the 18 latest models of four major task domains: Image Classification (IC), Image Segmentation (IS), Object Detection (OD), and Natural Language Processing (NLP).
+We evaluate the effect of GPU acceleration on the relative latency across GEMM and non-GEMM operators in collected models and conduct deep-dive analysis on the impact of different hardware platforms (mobile, workstation, and datacenter), deployment software, and common optimizations (operator fusion and quantization).
+Based on our case studies, we highlight that the non-GEMM performance challenge is common in accelerated inferences and existing optimization techniques (e.g., operator fusion) cannot completely address the challenges.
+Also, we demystify the most time-consuming non-GEMM operators in each model, which will help the research community identify non-GEMM operators to be optimized.
+To facilitate such research in non-GEMM-oriented optimization techniques, we build an open-source benchmark specialized in non-GEMM performance anlysis, NonGEMM Bench, which will be released after publication.
+NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch [46], ONNX [7], and TensorRT [43], in addition to the preset of non-GEMM operators collected from the selected 16 popular models, which provides desired flexibility to users for follow-up research.
+We summarize our contributions as follows:
   • We shed a light on the changed landscape of Amdhal’s law in ML system design, which shows the increased importance of non-GEMM operators in systems with GEMM accelerations.
-  • We propose NonGEMM Bench, which provides analysis on non-GEMM operators in the 18 latest popular models from Hugging Face and Torchvision.
-  • To facilitate non-GEMM operator-aware optimizations, we organize a microbenchmark of non-GEMM operators collected in the popular models and provide realistic input arguments collected from realistic input data.
-  • Using NonGEMM Bench, we perform case studies on three different hardware configurations ranging from mobile, workstation, and to data center, and show the non-GEMM operators are becoming major consideration across all platforms.
-  • We identify different set of dominant non-GEMM operators depending on the task domain and deployment SW flow, which motivates domain-oriented specialization and shows the importance of deployment software’s role for non-GEMM operators.
-2 Background
+  • We perform case studies on two different hardware configurations, workstation and data center, and show the non-GEMM operators are becoming a major consideration across all platforms.
+  • We identify different dominant non-GEMM operators depending on the model and deployment software flow, which indicates that non-GEMM optimization need to be specialized for each model and deployment software.
+  • We analyze the impact of a common non-GEMM-aware optimization, operator fusion and show that operator fusion does not completely mitigate non-GEMM bottleneck for all models, which motivates follow-up research in non-GEMM performance optimizations
+  • We evaluate the performance of non-GEMM operators with quantization and show the non-GEMM bottleneck aggravates with data.
+  • We will open-source NonGEMM Bench, an expandible benchmark flow that enables thorough non-GEMM performance characterization for any model supported by ONNX runtime, Tensor RT, and PyTorch, to facilitate non-GEMM-oriented research.
+II Background
 Figure 2: A description of Linear (a) and Conv1D (c) operators as GEMM operators example, and that of non-maximum suppression [21] (b) and Layer Normalization [6] (d) as a non-GEMM operator example.
-To introduce NonGEMM Bench, we discuss the ML operators and recent ML models.
-2.1 ML Operators
+II-A ML Operators
 ML operators are the building blocks for ML models, which define the computation over one or multiple input tensors.
-Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization functions, as listed in Figure 2.
-Operators can be categorized into two classes: GEMM operation-based and the others, which we term as non-GEMM operators.
+Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization, as listed in Figure 2.
+We categorize operators into two classes: GEMM operation-based and the others, which we term as non-GEMM operators.
 We discuss each class of ML operators next.
-2.1.1 GEMM-based Operators (GEMM Operators)
+II-A1 GEMM-based Operators (GEMM Operators)
 GEMM-based operators (or GEMM operators) refer to all the ML operators that can be represented as a matrix multiplication operation, which include linear, Conv2d, and batched-matrix multiplication (BMM).
@@
 However, they have regular computation patterns that can be summarized as a perfectly nested for loop.
-The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [33, 26, 46]
-With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [14]) and accelerators [23].
-2.1.2 Non-GEMM Operators
+The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [45, 33, 66] With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [13]) and accelerators [29].
+II-A2 Non-GEMM Operators
 Non-GEMM operators refer to all ML operators other than GEMM operators.
@@
 Because of such characteristics distinguished from GEMM operators, optimization methodologies for GEMM operators cannot be applied to accelerate non-GEMM operators.
-To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 18 recent models in the computer vision and natural language processing domains.
-We select models based on their popularity in the Hugging Face community to obtain realistic workload.
-We list the models we investigated in Table 1.
-Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table 2.
+To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 16 recent models in the computer vision and natural language processing domains.
+We select models based on their popularity in the Hugging Face to obtain realistic workload.
+We list the models we investigated in Table I.
+Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table II.
   • Normalization.
-Normalization operators regularize the data range across a selected dimension using the mean and standard deviation statistics.
-Examples include BatchNorm [22] and LayerNorm [7], which are widely adopted in computer vision and NLP ML models [18, 5].
-Such variants defer by the choice of the normalized dimension (e.g., BatchNorm normailizes a tensor across the batch dimension).
+Normalization operators regularize the data range across a selected dimension using the mean and standard deviation.
+Examples include BatchNorm [26] and LayerNorm [6], which are widely adopted in computer vision and NLP models [21, 4].
 RMS Norm [67], which is adopted in recent large language models [58], is another example of the normalization function.
@@
   • Memory Operators.
-Memory operators are responsible for the allocation and the layout modification of tensors inside the memory.
-Examples cover the reshape and concatenation (cat) operators.
-Reshape operators do not create new tensors.
-They modify the shape (e.g.
+Memory operators are responsible for the memory allocation and the layout modification of tensors.
+For example, Reshape modifies the shape (e.g.
 dimension order) of a tensor and return a new view of the tensor following the new dimension order.
-While reshape operators do not allocate additional memory, the cat operator allocates new memory to store a newly created tensors by concatenating the input tensors along a new dimension.
-  • Arithmetic.
-Arithmetic operators cover non-GEMM arithmetic operators applied on tensors in an element-wise or reduction pattern.
-Element-wise division applied to scale the elements of tensors in the attention block as shown in Figure 3 (c) is an example of such operators.
+  • Element-Wise Arithmetic.
+Elemennt-wise arithmetic operators cover non-GEMM operators applied on individual elements in a tensor.
+For example, Figure 3 (c) contains an element-wise division applied to scale the elements of tensors in the attention block
   • RoI Selection.
@@
 Figure 3: Architectures of three popular ML model families.
-2.2 ML Models and Popular Tasks
-The heterogeneity in non-GEMM operators enabled ML engineers and researchers to build models supporting a wide range of tasks (e.g.
+II-B ML Models and Popular Tasks
+The heterogeneity in non-GEMM operators enabled ML developers to build models supporting a wide range of tasks (e.g.
 Computer Vision and NLP).
@@
 Figure 4: An overview of NonGEMM Bench flows
-3 NonGEMM Bench
-Understanding the performance landscape of ML models with non-GEMM workloads necessitates a specialized benchmark.
-A good benchmark targeting non-GEMM operators should 1) capture the operator level performance breakdown not only in a standalone scenario, but also in an end to end inference to understand the effects of the non-GEMM operators.
-Moreover, it should analyze 2) frequently used models in 3) diverse task domains on 4) real data to reflect the actual performance of realistic application.
-ML Benchmarks available today (e.g., MLPerf [36]), unfortunately, do not satisfy the requirements since they do not focus on the non-GEMM operators.
+III Performance Characterization Methodology
+To understand the realistic performance landscape of the latest ML models with non-GEMM workloads, we must (1) capture operator level performances in end-to-end inferences to understand, (2) use widely-used models by the research community and industry, (3) cover diverse task domains, and (4) use real datasets.
+However, ML Benchmarks available today (e.g., MLPerf [52]), unfortunately, do not satisfy all the requirements since they do not focus on the non-GEMM operators.
 Long-tail bench [34] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domain.
-To better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench, that satisfies all the requirements for a non-GEMM operator specific benchmark.
-NonGEMM Bench provides operator-level breakdown of execution time at the computational/operator graph level in end-to-end inference scenarios, which enables detailed non-GEMM operator performance analysis (e.g., impact of each operator type on the end-to-end performance), as we present in Section 4.
-In addition, to facilitate non-GEMM operator-focused system optimizations, NonGEMM Bench includes a micro-benchmark of non-GEMM operators with concrete input arguments and tensors collected from realistic datasets and models.
-We discuss the datasets and models adopted in NonGEMM Bench and describe the structure of NonGEMM Bench in detail next.
-3.1 NonGEMM Bench Models
-To capture the performance in the latest ML workload, we select 18 highly downloaded models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench.
-Table 1 lists the NonGEMM Bench model registry (Figure 4) which contains 18 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench.
+Therefore, to facilitate our non-GEMM operators analysis and better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench.
+NonGEMM Bench provides operator-level breakdown of end-to-end inference latency in the operator graph level, which enables detailed non-GEMM operator performance analysis, as we present in Section IV.
+To capture the performance in the latest ML workload, we select 15 highly downloaded (more than 10K downloads on average) models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench and our analysis.
+We discuss the datasets and models in NonGEMM Bench in detail and describe the structure of NonGEMM Bench next.
+III-A Evaluated Models
+Table I lists the NonGEMM Bench model registry which contains 15 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench.
 The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
-3.1.1 Computer Vision (CV)
-Deep Learning is widely used in the CVdomain [25]  [19]  [38]. NonGEMM Bench covers three important tasks of the domain:
+III-A1 Computer Vision (CV)
+Deep learning is widely used in the CV domain [32]  [22]  [53]. NonGEMM Bench covers three important tasks in the domain:
 Image Classification (IC).
 Image classification refers to a CV task that identifies a class label of a given image.
-Image classification models extract features (i.e., high-level and -dimensional information of the input image) from an input image and reports the class label utilizing the features.
-NonGEMM Bench includes eight most popular IC models in HuggingFace [1], which includes ResNet50  [19], MobileNetv2  [39], three variants of Vision Transformer [25], and three variants of Swin Transformer  [29].
+Image classification models extract features (i.e., high-level and dimensional information of the input image) from an input image and report the class label utilizing the features.
+NonGEMM Bench includes eight most popular IC models in HuggingFace [1]: Three variants of Vision Transformer [32], and three variants of Swin Transformer  [37].
 Image Segmentation (IS).
 Image segmentation refers to a computer vision task that identifies the area in a image for each class.
-Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels belong to each object.
+Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels that belong to each object.
 NonGEMM Bench includes two state-of-the-art IS models: Segformer  [64] and MaskFormer  [10].
@@
 Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object.
-Like aforementioned models, OD models extract features.
-In addition to that, OD models generate region proposal, which refers to the candidate location and bounding boxes for objects in an image.
-Using Region of Interest (RoI) processing non-GEMM operators listed in Table 2, OD models refine the raw region proposal generated by a region proposal network.
-The refined RoIs are used as inputs to the classifier CNN at the end, and the classifier determines the class label of object inside each refined RoI.
+OD models extract features and generate region proposals, which refer to the candidate locations and bounding boxes of objects.
+Using Region of Interest (RoI) processing non-GEMM operators in Table II, OD models refine raw region proposals generated by a region proposal network.
+The refined RoIs are used as inputs to the CNN classifier at the end, and the classifier determines the class label of objects inside each refined RoI.
 NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [53], MaskRCNN  [21], and DETR  [9].
-3.1.2 Natural Language Processing (NLP)
+III-A2 Natural Language Processing (NLP)
 NLP refers to tasks involving the analysis and understanding of human (natural) language.
 NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [68] [8].
-Transformer [42] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state of the art generative large language models like GPT  [35] and Llama  [41].
+Transformer [59] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state-of-the-art generative large language models like GPT  [50] and Llama  [58].
 Figure 3 (c) shows the layer architecture of GPT’s transformer.
-It consists of a self attention block build by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table 2).
+It consists of a self-attention block built by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table II).
 NonGEMM Bench includes five popular NLP models  [3]: Bert [31], three variants of GPT2 [4], and Llama2-7b [58].
-3.2 Benchmark Suite
-NonGEMM Bench provides two flows: end-to-end inference and microbench flows.
-The End-to-end flow evaluates the performance of operators when the model is running inference on dataset data.
-The Microbench flow performs micro-benchmarking on the operator by running it in a standalone fashion using synthetic data with realistic dimensions.
-We discuss input, output, and the benchmark software structure for each flow.
-3.2.1 NonGEMM Bench Inputs
-NonGEMM Bench receives target deployment software flow and other configurations as inputs.
+III-B Profiling Methodology
+To enable the thorough analysis of the landscape of non-GEMM performance presented in Section IV, NonGEMM Bench provides an end-to-end profiling flow to breakdown the inference latency at the operator level.
+We discuss input, output, and the benchmark software structure.
+III-B1 NonGEMM Bench Inputs
+NonGEMM Bench receives target deployment software flow (PyTorch, ONNX runtime, or TensorRT), dataset, and other configurations listed in Figure 4 as inputs.
 Optionally, users can provide custom models with associated datasets to profile their workloads.
-By default, NonGEMM Bench runs the 18 models in Table 1.
-We describe each input in detail as follows:
+By default, NonGEMM Bench runs the 15 models in Table I.
+We describe each input in detail as follows.
 Models.
-As described in  Section 3.1, NonGEMM Bench offers a registry of pre-selected popular ML models for users to analyze.
+As described in  Section III-A, NonGEMM Bench includes a registry of 15 selected popular ML models.
 Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators.
 Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
-Deployment Flow. NonGEMM Bench deploys its models using two popular inference frameworks: ONNX Runtime [16] and PyTorch [34] eager mode, TorchScript, and TorchDynamo introduced in PyTorch 2.0 [6].
+Deployment Flow. NonGEMM Bench deploys its models using three popular inference frameworks: ONNX Runtime [17], PyTorch [46] TensorRT [43].
 Datasets.
@@
 We use ImageNet 2012  [15] and MS COCO  [36] for computer vision tasks.
-As for language models, we use wikitext dataset [31] available on HuggingFace [4].
-Misc. Configurations. This includes all other configurations for profiling runs, which includes the batch size, the number of profiling iterations, and the target device.
-Table 1: Tasks and Models in NonGEMM Bench
+As for language models, we use wikitext dataset [39] available on HuggingFace.
+For custom models, users can specify their own dataset for the model.
+Misc. Configurations. This includes the batch size, the number of profiling iterations, and the target device.
+TABLE I: Tasks and Models Evaluated in Section IV
 Application Models # Params Dataset
-Image Classification ResNet50 [19] 0.85M ImageNet [15]
-MobileNetv2 [39] 3.4M
-ViT large 16 [25] 307M
-ViT huge 14 [25] 632M
-Swin tiny [29] 29M
-Swin small [29] 50M
-Swin base [29] 88M
-Object Detection FasterRCNN [38] 42M COCO [28]
-MaskRCNN [18] 44M
+Image Classification (IC) ViT base (Vt-b) [32] 307M ImageNet  [15]
+ViT large (Vt-l) [32] 307M
+ViT huge (Vt-h) [32] 632M
+Swin tiny (Sw-t) [37] 29M
+Swin small (Sw-s) [37] 50M
+Swin base (Sw-b) [37] 88M
+Object Detection (OD) FasterRCNN (FRCNN) [53] 42M COCO  [36]
+MaskRCNN (MRCNN) [21] 44M
 DETR [9] 41M
-Segmentation Maskformer [12] 102M COCO [28]
-SegFormer [45] 3.7M
-Language Models GPT2 [5] 117M wikitext [31]
-GPT2 Large [5] 762M
-GPT2 X-Large [5] 1.5B
+Image Segmentation (IS) Maskformer (MF) [10] 102M COCO  [36]
+SegFormer (Seg) [64] 3.7M
+Natural Language Processing (NLP) GPT2 [4] 117M wikitext  [39]
+GPT2 Large (gpt2-l) [4] 762M
+GPT2 X-Large (gpt2-xl) [4] 1.5B
 Llama 2-7B [58] 7B
 Bert [31] 110M
-3.2.2 End-to-End Inference Flow
-The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset, which consists of two stages.
-The first stage extracts the operator graph of a given model and preprocess input data.
-The second stage records the latency and energy of the inference runs.
-Graph Extractor functions convert the model to be evaluated into a computational graph exposing the GEMM and non-GEMM operators for profiling.
-Based on the selected deployment flow, the graph extractor calls the corresponding APIs to trace through the model and generate the operator graph.
-For the ONNX Runtime deployment flow, the NonGEMM Bench harness calls ONNX graph exporters from the HuggingFace transformers library [44] or from the PyTorch libraries.
-For the PyTorch deployment flow, the graph extractor relies on the torch.fx [37] Tracer to obtain the computational graph.
-The Tracer propagates a proxy data structure into the model and captures the operators as graph nodes.
-Data Preprocessing functions are model specific functions that feed the profiling functions with input tensors.
-First, they fetch the raw data from the specified dataset and perform data cleaning.
-Empty sequences are removed from text data and images are re-scaled to each model’s specific input resolution.
-Second, they apply tokenization or image to tensor transformation to generate the input tensors.
-End-to-End Inference Profiling Functions receive operator graphs and input tensors from the harness functions and run profiling iterations.
-The latency profiling functions record the operator breakdown of the end to end latency as well as the shape of the operators input tensors.
-Similar to the graph extractors, the latency profiling functions are deployment flow specific.
+III-B2 End-to-End Inference Flow
+The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset.
+First, the Graph Extractor extracts the operator graph of a given model based on the selected deployment flow.
+NonGEMM Bench calls graph exporters from the HuggingFace transformers library [62] or from the PyTorch libraries.
+Second, model specific Data Preprocessing functions fetch the raw data from the specified dataset, clean the data, and apply tokenization or image to tensor transformations to generate input tensors for the profiling stage.
+Third, End-to-End Inference Profiling functions run the profiling iterations on the received the model and its inputs.
+The profiling functions are deployment flow specific.
 For ONNX Runtime, we invoke Execution Provider profiling in the inference session.
-Currently, we only support operator level profiling for CPU and CUDA execution providers because profiling support to other ONNX RUNTIME backends is not stable yet.
-In PyTorch eager mode, we modify the torch.fx Interpreter to invoke the PyTorch profiler at each node of the operator graph.
-In TorchScript and TorchDynamo, we currently rely on the PyTorch Profiler.
-The energy profiling functions record the end to end inference power and energy consumption.
-We utilize system specific profiling tools to obtain CPU and GPU power profiles: NVIDIA’s System Management Interface (nvidia-smi) and AMD micro-Prof ($\mu$Prof) software profiling analysis tool.
-3.2.3 MicroBench Flow
-The MicroBench flow focuses on individual non-GEMM operators, which allows users perform deep analysis of their systems targeting non-GEMM operators.
-The flow consists of MicroBench Extractor, the NonGEMM Bench Operators Registry, and the MicroBench Operator Profiling (shown as ”u-Bench op Profiling in Figure 4).
-MicroBench Extractor.
-This component populates the non-GEMM operators captured during end-to-end flow of NonGEMM Bench to the microbench operator registry.
-In addition to the operators, the registry stores the shape of input tensors to the non-GEMM operators as well as the implementation in their parent model.
-Such a step is crucial for keeping the microbenchmark realistic.
-MicroBench Operator Profiling function (”u-Bench op Profiling” in Figure 4) fetch the non-GEMM operators from the registry and profile their performance utilizing latency/energy profiling in the end-to-end flow.
-For each operator, it generates input tensors using the recorded shapes and proceeds to evaluate the latency and the energy of the operator.
-3.2.4 NonGEMM Bench Outputs
-NonGEMM Bench generates many statistics orgainzed into three categories: performance/cost, workload, and non-GEMM-specific reports.
-The performance/cost report includes key performance metrics such as the end-to-end latency with operator level break-downs, energy, and peak memory usage.
-The workload report includes the shape of the tensors captured during inference on realistic data.
-Finally, the non-GEMM report provides insights on non-GEMM operators such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework), non-GEMM operator trace on different domains, and so on.
-Table 2: Non-GEMM operators in eight popular model variants selected from Hugging Face listed in Table 1.
+Currently, we only support operator level profiling for CPU and CUDA execution providers [17] because profiling support to other ONNX RUNTIME backends is not stable yet.
+For PyTorch, we use the PyTorch Profiler [49].
+For TensorRT, we leverage the TensorRT open source software framework profiling APIs [44].
+III-B3 NonGEMM Bench Outputs
+NonGEMM Bench generates many statistics organized into three categories: performance, workload, and non-GEMM-specific reports.
+The performance report includes key performance metrics such as the end-to-end latency with operator level break-downs and peak memory usage.
+The workload report includes the types of operators and the shape of the tensors for each operator captured during inference on realistic data.
+Finally, the non-GEMM report provides insights on non-GEMM operators, such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework) and non-GEMM operator trace on different domains.
+TABLE II: Non-GEMM operators in eight selected model variants from Table I and their characteristics.
 Operator Group Operator Model Single Operation Single Operand Non Linearity Dynamicity Reduction Example Input Shape
@@
 FrozenBatchNorm2d MaskRCNN [1, 1024, 50,68]
-FrozenBatchNorm2d DETR [2, 850, 256]
+FrozenBatchNorm2d DETR [1, 2048, 25, 34]
 LayerNorm DETR [2, 850, 256]
-Arithmetic Add Segformer [2, 16384, 32]
+Elmt-wise Arithmetic Add Segformer [2, 16384, 32]
 Mul Llama-2 [1, 10, 11008]
@@
 Interpolation Interpolate Segformer [2, 256, 128, 128]
-4 Case Studies
-We perform a thorough analysis of the performance data generated by running NonGEMM Bench on multiple hardware platforms show in  Table 3.
-We collect the data by running NonGEMM Bench in PyTorch eager mode on three systems configuration ranging from mobile CPUs to data center grade CPUs and GPUs to capture the behavior of the operators when subjected to various acceleration capabilities.
-We notice that for most of the models, the time spent to compute non-GEMM operators increases noticeably with respect to the total execution time when running the model on the GPU.
-Table 3: Hardware platform configurations used for NonGEMM Bench case studies.
-Class CPU GPU
-Data Center AMD EPYC 7763 NVIDIA A100
-Workstation Intel i9-13900K NVIDIA RTX 4090
-Mobile Intel i7-13700HU NVIDIA RTX 4060m
-4.1 Impact of Non-GEMM on Each Task Domain
-We discuss how non-GEMM operator’s contribution to the end-to-end inference and energy changes depending on task domains.
-4.1.1 Image Classification
-As shown in Figure 5, we observe a clear change in relative break-down latency across operators when using a GPU.
-For example, the normalization operation is minor before GPU acceleration, but it becomes the dominant operator with a GPU, accounts for over 50% of the total execution time, on average.
-We observe similar patterns for all the other cases: GEMM operators is dominant on CPU-only systems, but non-GEMM operators become significant or considerable after GPU acceleration.
-For the batch size of 1, the relative execution time of Non-GEMM operators for running ViT-b16 and ViT-l16 increases from 23.5% and 18% when running on a CPU only platform to 60% and 55%, respectively, on the GPU as shown in  Figure 5.
-Similarly, for the workstation GPU system, the non-GEMM operators execution time proportion increases to 56% and 45% for ViT-b16 and ViT-l16 to almost split the total execution time evenly with GEMM operators.
-For the three Swin transformers with a batch size of 1, the non-GEMM operators contribution to the overall latency increases as well when running the models on the GPU.
-The small Swin transformer model spends 55% of the computation time executing non-GEMM operators on the data center GPU system compared to a modest 33% contribution on the CPU only system.
-On the workstation GPU system, the small Swin transformer also computes non-GEMM operators for 57% of the time, a 2x increase compared to the workstation CPU system.
-Impact of the batch size.
-When increasing the batch size, the latency distribution across GEMM and non-GEMM operators of IC models running on GPU follows the same trend as that of single batch.
-However, this trend is less noticeable for the largest vision transformer, ViT huge (632M parameters).
-For a batch size of 8, the non-GEMM operator latency is 43% and 51% for ViT-l16 and base Swin transformer respectively (Figure 5) running on the data center GPU, an 3.9$\times$ and 3$\times$ increase compared to the CPU only results.
-However, the huge vision transformer only shows a 1.25x increase in the latency proportion of non-GEMM operators when accelerating with a GPU.
-The non-GEMM operators with the most significance to the execution time of classification models are presented in Figure 5.
-As shown in  Table 4, Normalization operators are the most expensive non-GEMM operator in classification models accounting for an average of 17% of the execution time on the data center GPU, followed by Arithmetic operators with 14% (Figure 5).
-Based on NonGEMM Bench collected data, image classification models require  48% of execution time on non-GEMM operators on GPU systems, on average, across all the evaluated models.
-4.1.2 Object Detection
-Non-GEMM operator’s latency is also noticeable for object detection with a GPU.
-On the data center, non-GEMM operators dominated the execution time of FasterRCNN by consuming 82% of the total inference time when we accelerate the workload compared to 46% when running the CPU only (Figure 5).
-DETR and MaskRCNN showed a similar behavior after acceleration, with a 2x increase in the time spend on non-GEMM operators.
-DETR computed non-GEMM operators for 77% of the inference time on the data center GPU, while MaskRCNN spent 80% on the data center and workstation GPUs.
-For object detection models, Normalization operators exhibit the highest execution time followed by activation and memory operators as indicated in (Figure 5).
-All detection models in NonGEMM Bench use normalization functions with custom implementations that do not leverage available library normalization kernels.
-This leads to additional overhead to execute the normalization function.
-Overall, the evaluation data shows that detection models get dominated by non-GEMM operators (78%) when executing the workloads on a GPU system.
-4.1.3 Segmentation
-For segmentation models, the non-GEMM operators account for 47% of the execution time when running the inference on the CPU on average.
-Upon running the model on the GPU, this number slightly increases to 53%.
-Figure 5 shows memory and arithmetic operators to be the most latency expensive non-GEMM operator on GPU accounting for 17% and 12.5% respectively.
-For the SegFormer model, Normalization operators are the most expensive (15% of the execution time), while Memory operators dominate the non-GEMM operators of Maskformer (24%).
-4.1.4 Language Models
-GPT2 and Bert executed non-GEMM operators for 77% and 57% of the time when running on a single batch as indicated in Figure 5.
-For GPT2-large, the non-GEMM operators clearly dominated the execution time on GPU systems as well.
-Across the three GPU HW classes, non-GEMM operators occupy 76% of execution time on average.
-Table 4 show that Activation (23% on GPT2) and element wise arithmetic (20% on Bert) operators are the most expensive non-GEMM operators in language models running on GPU.
-Table 4: NonGEMM Bench Most Expensive NonGEMM Operator Group for selected NonGEMM Bench Models Running on Data Center GPU Platform.
-Models Model Alias Batch Size Operator Group % of Exec. Time
-ViT Base 16 Vt-b 1 Norm 20
-8 Norm 17
-ViT Large 16 Vt-l 1 Norm 20.5
-8 Arith 17
-Swin Tiny Sw-t 1 Norm 19
-8 Norm 18.5
-Swin Small Sw-s 1 Norm 18
-8 Norm 17
-Swin Base Sw-b 1 Norm 18
-8 Norm 18
-FasterRCNN FRCNN 1 Norm 60.5
-2 Norm 59
-8 Norm 49
-MaskRCNN MRCNN 1 Norm 58
-2 Norm 56
-8 Norm 49
-DETR DETR 2 Norm 40
-GPT2 gpt2 1 Act 23
-64 Act 16
-GPT2 X-Large GPT2-XL 1 Act 23
-64 Act 9
-Llama2-7B Llama2 1 Arith 23
-Bert bert 1 Arith 20
-64 Arith 10
+IV Case Studies
+We perform a thorough performance analysis of models in Table I on multiple hardware platforms listed in  Table III.
+The evaluated hardware platforms span workstation and data center-class CPUs and GPUs.
+We employ PyTorch, ONNX runtime, and TensorRT as deployment flows to investigate their impact on the performance horizon.
+Using the aforementioned workload, hardware platforms, and deployment flow, we characterize the GEMM/non-GEMM performance for each model in NonGEMM Bench and investigate the impact of the hardware platform and deployment flow.
+In addition, we also investigate the impact of widely adopted optimizations (operator fusion and quantization) and show that the non-GEMM performance is an issue even after applying common optimizations.
+TABLE III: Hardware platform configurations used for case studies.
+ID Class CPU GPU
+Device Device Mem. TOPS
+A Datacenter AMD EPYC 7763 Nvidia A100 80 GB 624
+B Workstation Intel i9-13900K Nvidia RTX 4090 24 GB 660
+IV-A Non-GEMM Performance Characterization Results
+We first conduct a performance characterization study using PyTorch and present the characterization results on datacenter and workstation platforms (Platforms A and B in Table III) in Figure 5.
+We focus on those two platforms since some models (e.g, large language models) do not fit into the mobile devices.
+As expected, the relative contribution of non-GEMM operators to the end-to-end latency significantly increases after GPU acceleration, on average, from 17.2% to 42.3%.
+However, we observe each model show different trends, mainly affected by the non-GEMM operator types and the number of GEMM and non-GEMM operators in the original model.
+We summarize the most time-consuming non-GEMM operators in Table IV from the datacenter class platform (Platform A), which shows the diversity of the dominating non-GEMM operators in each model.
+We highlight some notable models in each task category and delve into the details of their non-GEMM performance.
+IC Task: Swin Transformer.
+For every Swin Transformer variant(Sw-t, -s, and -b), the memory operator group is the most latency-dominant non-GEMM operator group, which accounts for 32% of the total latency, on average, on data center platform with GPU acceleration.
+Those memory operators originate from the Swin Transformer’s unique window shape (cross-shaped) [37], which is not well-aligned with memory layout of tensor data organized in dimension orders.
+OD Task: DETR.
+After GPU acceleration, DETR has shown significant non-GEMM operator presence, which accounts for 65.8% of the total latency, on average.
+The major source of the non-GEMM latency is in the normalization operators, whose percentages are 35% and 32% on Platforms A and B, respectively.
+We observe the normalization functions are based on a custom implementation, which are identified as independent kernel.
+Kernel launch overheads accumulated for independent runs for the custom normalization function significantly contribute to the non-GEMM latency.
+However, we also observe that an advanced deployment software (TensorRT) can fuse those operators and significantly improve the non-GEMM performance.
+We discuss the details later in Section IV-B.
+IS Task: Maskformer.
+MaskFormer utilizes Swin Transformer as its backbone, which introduces many memory operators as we discussed for Swin Transformer.
+As a result, memory operator is the dominant non-GEMM operator, which accounts for 40.8% of the total latency, on average, as we can observe in Figure 5 (f) and (n).
+NLP Task: GPT2.
+The latencies of non-GEMM operators in GPT2 variants are considerable, which account for 45.0%, on average, as we observe in Figure 5 (h) and (p).
+The dominant non-GEMM operator is activation, GELU, which required 26.4% of the total latency.
+TABLE IV: Most time-consuming non-GEMM operator groups for selected Models (Platform A, with GPU acceleration, average across batch sizes).
+Task Domain Model Operator Group Latency Percentage (%)
+Image Classification Vt-b Norm 14.0
+Vt-l Norm 13.3
+Vt-h Norm 11.2
+Sw-t Memory 31.8
+Sw-s Memory 33.1
+Sw-b Memory 32.8
+Object Detection FRCNN Elmt-wise Arith. 34.4
+MRCNN Elmt-wise Arith. 33.6
+DETR Norm 34.8
+Image Segmenation MF Memory 40.8
+Seg Normalization 17.4
+NLP gpt2 Act 30.2
+GPT2-L Act 29.9
+GPT2-XL Act 28.1
+Llama2 Norm 14.9
+bert Norm 13.1
+Summary.
+We observe GPU acceleration significantly increase the percentage of non-GEMM operators and increase the importance of non-GEMM operators in performance optimization.
+Also, the dominant operators are diverse in each model, which indicates an optimization technique tailored for a single operator cannot fully address the non-GEMM performance problem.
 Figure 5: Breaking down the execution time of models running on the Data Center (CPU only) and (CPU + GPU) configurations across operator groups.
-Figure 6: Breaking down the execution time of models running on the Workstation (CPU only) and (CPU + GPU) configurations across operator groups.
-Large language model’s latency is dominated by GEMMs when running on Data center, consisting of 79% of the total inference time on average across platforms.
-When utilizing the GPU, the GEMM latency no longer bottlenecks the total execution time.
-The GEMM operators time drops to 30% and 22% for Llama2-7b and GPT2-XLarge respectively with non-GEMM operators executing for 74% on the GPU configuration.
-Figure 7 (a) suggests that the bottleneck are activation, arithmetic, and memory operators.
-Element-wise arithmetic operators are the most time-consuming non-GEMM operators for Llama2-7b, accounting for 23% of the total runtime (Table 4, while activation functions occupy 23% of GPT2-XL total execution time.
-GPT2 models implement their own GELU activation in the Hugging Face library which does not have a direct mapping to a GELU GPU kernel when running in Pytorch eager mode.
-This leads to additional overheads due to running multiple kernels to compute the activation.
-Similarly, Llama uses a normalization function that does not map to a GPU kernel in eager PyTorch which leads to similar overheads that increase the latency of the non-GEMM operators.
-4.2 Impact of Deployment Flow on Non-GEMM
-Running in eager PyTorch, Non-GEMM operators occupy large proportions of the execution time for various applications running on the GPUs , as summarized in Table 4.
-Nevertheless, deployment frameworks (e.g.
-[16, 11]) are typically used to improve the performance beyond the eager mode [6].
-They apply transforms to the computational graph level and leverage accelerators (if available) to optimize the inference runs.
-To understand how deployment frameworks affect inference performance distribution across operators, we evaluate NonGEMM Bench models with ONNX Runtime (ORT).
-ORT is a cross platform ML framework that accelerates ML model inferencing [16].
-It leverages hardware accelerators (e.g.
-GPUs) and operator graph level optimizations to improve the performance of ML models.
-We run selected models with the CUDA execution provider and observe the following results.:
-Non-GEMM operators account for a significantly larger proportion of the total execution latency (Figure 8) when running on different grade GPUs.
-This is aligned with the trend observed with the PyTorch eager execution on GPU.
-On average, the time spend on non-GEMM operators increases from 52% to a 73% after deploying the models on GPU with the ORT cuda execution provider.
-Memory operators almost dominate the execution time in three application domains, classification, detection and language models by occupying 66%, 44%, 59 % respectively.
-For instance, memory operators are the most latency expensive non-GEMM operators in GPT2-XLarge and Llama2-7b as they consume 76% and 59% of the total inference time respectively (Figure 7).
-In detection models, memory operators consume 37% of the total inference time in ORT as well (Figure 8).
-One reason leading memory operators to dominate the inference latency is their high frequency, notably in the large language models.
-They consist of 80% and 62% of the total operators in the GPT2-XLarge and Llama2-7b .
-Consequently, the cumulative latency to run these operators becomes the inference latency bottleneck.
-Another reason is the limitations in operator support for GPU deployment in ORT.
-Memory operators are a subset of non-GEMM operators that the cuda execution provider does not fully support, hence it assigns them to be computed on the CPU.
-In the case we hit an unsupported operator while executing on the GPU, the runtime falls back to CPU to carry the computation.
-As a result, additional overheads to transfer the data between devices are added to the operators execution time that increases the latency of the fallen-back operators, which aggravates frequently used operators.
-Despite dominating the execution time in ORT, memory operators were not the main non-GEMM bottlenecks PyTorch eager mode for the evaluated models.
-Figure 7 highlights the differences in operator latency distribution for GPT2-XLarge and Llama2-7b running in PyTorch eager mode and ORT.
-This indicates that the deployment flows have distinct effects on the execution of non-GEMM operators in an end to end inference.
-Figure 7: The impact of deployment software toolchain into the latency breakdown. (a) PyTorch (Eager Mode) and (2) ONNX Runtime on a data center class GPU (A100).
-Figure 8: The latency breakdown of the inference using ORT on two platform configurations: mobile and Data Center.
-4.3 Key Observations and Insights
+IV-B The Impact of Deployment Flow on Non-GEMM Performance
+Deployment flows such as ONNX runtime [17] and TensorRT [43] are widely used for serving model inferences.
+Such flows apply various optimizations to each model, which includes the computational graph level optimizations (e.g., operator fusion) and backend assignment (e.g., utilizing Tensor Core in Nvidia GPUs).
+To understand the effect of deployment frameworks on the operator performance, we conduct two case studies (1) comparing PyTorch and ONNX Runtime (ORT) results on LLMs(focus: general optimizations w/o operator fusion) on LLMs, and (2) comparing PyTorch and TensorRT results (focus: operator fusion).
+IV-B1 Case Study 1: Non-GEMM Performance on LLMs across PyTorch and ORT
+We run two LLMs (GPT2 and Llama2) using the CUDA execution provider in ORT and compare their non-GEMM performance percentage in Figure 6.
+In the results, we observe the presence of non-GEMM operators significantly increase, from 52.6% to a 80.75%, on average.
+We observe the percentage of memory operators significantly amplifies after applying ORT, from 3.2% to 66.8%.
+Such a result originates from the limitation in operator support in ORT.
+Many memory operators in the evaluated LLMs are not supported by the cuda execution provider in ORT, which leads to inefficient execution on CPUs involving data transfer between a CPU and a GPU.
+Combined with the high frequency of those operators, the inference latency of memory operators increases significantly, as shown in Figure 6 (b).
+The results imply two insights: (1) Model deployment flow can significantly aggravate the non-GEMM performance challenge and (2) the dominant non-GEMM operators differ depending on the operator support in a deployment flow.
+Figure 6: The impact of deployment software toolchain into the latency breakdown on language models. (a) PyTorch and (2) ONNX Runtime on a data center class GPU (A100).
+Figure 7: The latency of non-GEMM operator in (a) Swin-b, (b) DETR, and (c) Segformer is still considerable despite applying operator fusion in TensorRT.
+The numbers at the top of each bar represent the latencies in milliseconds.
+IV-B2 Case Study 2: Non-GEMM Performance with Operator Fusion
+Operator fusion is a key optimization technique for accelerating inference workloads [14, 25, 57, 42, 43, 30].
+Fusion combines multiple operators in a single kernel to reduce the number of costly kernel launches and minimize the number of redundant offchip memory accesses [42].
+TensorRT [43] is a widely adopted inference framework released by Nvidia that applies the operator fusion technique targeting GPUs.
+Operator fusion in TensorRT detects specific patterns (e.g., three consecutive element-wise operators [43]) in the operator graph and fuses nodes captured in the patterns to enhance inference performance by reducing redundant memory accesses around non-GEMM operators.
+Impact of Operator Fusion.
+To understand the impact of operator fusion on non-GEMM performance horizon, we conduct a case study on four models, which compares the TensorRT (with fusion) and PyTorch (without fusion).
+We present the results in Figure 7, which shows the inference latency breakdown between GEMM and non-GEMM operators.
+We notice that fusion does not completely address the non-GEMM bottleneck in Swin-b model.
+For example, the percentage of non-GEMM operators in the total latency changes from 56.2% to 32.2%, on average, after operator fusion by TensorRT.
+The reduction in the percentage is based on the improvement of non-GEMM performance via operator fusion, which reduces 88.6% of latency after fusion, as summarized in Table V.
+However, the non-GEMM operators still account for 32.2% of total latency, which is considerable.
+This shows that operator fusion cannot eliminate the non-GEMM performance challenge and motivates further studies toward non-GEMM performance.
+Factors Affecting the Effectiveness of Fusion.
+Although most results indicate considerable impact of non-GEMM even after operator fusion, we observe operator fusion on DETR model is exceptionally effective.
+Therefore, we conduct a deep-dive study, investigating the percentage of fused non-GEMM operators (i.e., fusion rate) and performance improvements after fusion, as listed in Table V.
+We observe the strong non-GEMM performance improvements for DETR originates from high fusion rate of 30%, which led to 13.5$\times$ non-GEMM speedup.
+This leads to the large percentage reduction of non-GEMM in the total latency, from 67% to 18.5%, on average.
+However, the fusion rate is not the only factor that determines the non-GEMM speedup.
+For example, DETR and Segformer have similar fusion rates (30% andn 27%, respectively), but the amount of non-GEMM performance improvements are significantly different: 13.5$\times$ and 2.39$\times$, respectively.
+We analyze the execution trace and find that batch normalization is the main source of the differences.
+The main difference is the fusion pattern: most batch normalization operators (100% of total) in DETR were fused with GEMM-operators (CONV+BN+ReLu pattern) while those in Segformer were fused with other non-GEMM operators (97.8% of total).
+The results indicate that the effectiveness of operator fusion relies on not only the overall fusion rate but also the fusion patterns.
+Our observation confirms that the operator fusion cannot fully address the non-GEMM performance challenge, even if it can be very effective on some patterns.
+IV-C Non-GEMM Performance with Quantization
+Quantization refers to the model optimization technique, which reduces the bit precision of model weights and/or activations to enhance the computational performance and efficiency of DNN inference.
+Quantization is a widely-adopted technique [28, 40] including heavy models like LLMs [19, 35, 63, 16].
+For example, LLM.int8() [16] introduced an 8-bit - 16-bit mixed precision quantization scheme for LLMs.
+LLM.int8() is a state-of-the-art quantization method, which quantizes more than 99% of the linear layers in OPT LLM to an 8-bit precision.
+Therefore, we adopt LLM.int8() and characterize GEMM and non-GEMM performance to understand the impact of quantization on the non-GEMM operator performance problem.
+We conduct a case study on quantization using LLM.int8() on Llama3 LLM on the datacenter platform with GPU acceleration and present the results in Figure 8.
+In the results, we observe that non-GEMM operators dominate the latency after quantization, changing from 29.3% to 76.7%, on average.
+Such a significant shift in the latency distribution is mainly based on the GEMM performance improvements from 8-bit arithmetic, which reduced the latency by 38.2%, on average.
+However, based on our analysis on the execution traces, non-GEMM performance aggravates because the 8-bit data need to be dequantized and re-quantized for non-GEMM operators, which requires 16-bit floating point arithmetic.
+This led to 6510 additional non-GEMM operators into the computational graph, which led to a significant increase of the non-GEMM latency by 5.6$\times$ after quantization.
+Combined together, the overall percentage of non-GEMM operators in the total latency dominate after quantization, which makes non-GEMM operators as the major optimization target.
+In the case study on Llama3 8B, we observe the element-wise arithmetic operators originate from dequantization/requantization (DQRQ) process dominate in the inference latency, which adds 20% extra non-GEMM operators to the original computational graph.
+Also, we observe longer sequence length leads to higher percentage in the element-wise arithmetic operators.
+For example, as we increase the sequence length from 512 to 8192, the latency percentage of element-wise arithmetic operators increase from 31.8% to 63.8%.
+As current trends in the LLMs are toward longer sequence lengths [18, 5], the non-GEMM performance issues in longer sequences originating from DQRQ costs will aggravate, which motivates efforts in non-GEMM performance optimization.
+Figure 8: Non-GEMM operators latency contribution to the inference of a quantized Llama3 8B using 8 bit quantization [16]. The numbers at the top of each bar represent the latencies in milliseconds.
+TABLE V: The non-GEMM latency before and after applying fusion with TensorRT. The values between brackets represent the percentage w.r.t the total inference latency.
+Model Non-GEMM Non-GEMM Latency
+Fusion Rate Before Fusion After Fusion
+Swin-t 8.75% 7.53 ms (56.4%) 0.97 ms (39%)
+Swin-b 7% 14.59 ms (58%) 1.65 ms (32.25%)
+DETR 30% 32.17 ms (66.4%) 2.38 ms (18.5%)
+Segformer 27% 5.57 ms (41%) 2.33 ms (41%)
+IV-D Key Observations and Insights
 We summarize our main observations and insights:
-  • Accelerating GEMM operators using GPUs significantly shifts the performance horizon, which makes the non-GEMM operators main bottleneck of the systems: On average, non-GEMM breakdown increases from 27% (CPU-only) to 55% (CPU + GPU).
-  • We observe the dominance of non-GEMM operators after GPU acceleration across all the platforms (mobile, workstation, and datacenter) and task domains.
-  • The choice of deployment software flow affects the most time-consuming non-GEMM operators: In ORT, memory operators dominate the execution time across all models (on average, 56%), while PyTorch doesn’t show clear patterns.
-  • Dominant non-GEMM operator class vary depending on the task category: In IC models, normalization operators dominate (on average, 18.4% of the total execution time), while activation and element-wise arithmetic operators dominate in language models (17.75% and 17.6% for each, on average).
-  • The profiling results indicate that the non-GEMM operators are now the prime optimization target after GEMM acceleration.
-5 Related Works
-Table 5: Comparing NonGEMM Bench with existing benchmarks: MLPerf [36], LongTail Bench [27], TorchBench [17], and NonGEMM Bench (This work).
-Benchamrk MLPerf LongTailBench TorchBench NonGEMMBench (This work)
-Real Usage Driven
-Non-GEMM Focused
-Real Dataset Driven
-Plug Model & Profile
-Based on the characteristics of non-GEMM operators, and on the results shown in  Section 4, we clarify different features of key related works comparing against this work in Table 5.
-TorchBench [17] presents a comprehensive benchmark suite for the Pytorch framework consisting of 84 models whose selection was driven by real usage like NonGEMM Bench.
-MLPerf [36] is an industry standard state-of-the-art inference benchmark, which focuses on the end-to-end performance.
-LongTail Bench [27] is a micro-benchmark of NonGEMM operators.
-However, LongTail bench is a microbenchmark that only focuses on custom implementations of 34 operations, which cannot accept new operators nor profile various implementation of the same operator.
-As summarized in Table 5, all work shed a light to individual important aspect, but only NonGEMM Bench provides all the desired features for a non-GEMM operator-focused benchmark.
-6 Conclusion
-Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law in favor of non-GEMM operators.
-To guide future optimization, we need to understand the performance of non-GEMM operators in recent popular models.
-In this work, we analyze the performance non-GEMM operators in recent popular machine learning models.
-We construct NonGEMM Bench to study of non-GEMM operators in accelerated popular models reflecting realistic use scenarios and evaluate it on different hardware platforms.
-We show that the dominance of non-GEMM operators is occurring in all platforms with GPUs, and the deployment software’s impact is large.
-Our findings strongly motivate ML system researchers to consider non-GEMM operators as a major consideration.
+  • After GEMM acceleration, non-GEMM becomes a major optimization target regardless of the hardware platforms.
+  • Specialized optimization for one non-GEMM is not effective due to the diversity in dominant non-GEMM.
+  • Operator fusion cannnot fully address non-GEMM performance challenge: Although it can significantly improve non-GEMM performance, but its effectiveness heavily depends on the model.
+  • Operator support in deployment flows significantly affects the non-GEMM performance horizon.
+  • Quantization significantly aggravates the non-GEMM performance challenge due to the imbalanced speedup across GEMM and non-GEMM and quantization/dequantization costs around non-GEMM operators
+V Related Works
+Previous works characterized ML workloads and non-GEMM operators.
+However, the main limitations are listed below: State-of-the-art inference benchmarks  [20, 52] do not thoroughly analyze the performance of Non-GEMM operators.
+Other works characterizing non-GEMM operators focus on a specific application domain [48, 47, 65, 34] for their non-GEMM study.
+In [48],  [27], and  [47] the authors analyze the performance of non-GEMM operators, however they focus on a specific class of transformer language models.
+Similarly, [34] only studies a limited number of operators from computer vision models in a microbenchmarking flow.
+Previous works that also studied non-GEMM operator performance bottlenecks [12, 56, 60, 51], mostly focused on softmax operators in transformers for language models.
+Our work extends on the pioneering efforts of these works by studying non-GEMM operators from 15 models covering 4 various task domains.
+VI Conclusion
+Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law to have more presence of non-GEMM operators in the performance horizon.
+In this paper, we conducted a thorough performance analysis of non-GEMM operators in the latest models in various task domains and platforms.
+The results confirm the increasing importance of non-GEMM performance and show model optimizations we apply (e.g., quantization) can significantly aggravate the non-GEMM performance challenge.
+The dominant non-GEMM operators are diverse across models, which indicates that a specialized optimization targeting a specific operator cannot solve the non-GEMM performance challenge.
+We also show that non-GEMM-oriented optimization such as operator fusion cannot fully address the non-GEMM performance challenge.
+Our performance anlysis results imply that now we need to consider non-GEMM operators as a major optimization target and develop new hardware and software techniques to optimize non-GEMM performance.
+The non-GEMM profiling software we used in this study, NonGEMM Bench, will be released as open-source software, which will contribute to such follow-up research.

### v5

@@
 This led to advanced GPUs and accelerators available today, which provided significant boost in the GEMM performance compared to CPUs, aligned with the lesson from Amdahl’s law.
-However, accelerating GEMM has significantly shifted the Amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is not dominant.
+However, accelerating GEMM has significantly shifted the Amdahl’s law’s landscape for ML inference; due to the decreased GEMM execution time, the relative execution time of non-GEMM operators is now significant.
 Although the importance of non-GEMM performance is increasing, we have little knowledge about the non-GEMM performance horizon in the latest hardware platforms and models.
-Therefore, to guide non-GEMM-oriented optimizations, we conduct a thorough performance analysis of 16 widely adopted ML models in Hugging Face and Torchvision on workstation and data center platforms with/without GPUs.
+Therefore, to guide non-GEMM-oriented optimizations, we conduct a thorough performance analysis of 17 widely adopted ML models in Hugging Face and Torchvision on workstation and data center platforms with/without GPUs.
 We discover that non-GEMM performance bottleneck is a considerable issue across all the platforms and models, accounting for 11.3% to 73.6% of total latency, on average.
@@
 The GEMM-based operators dominate in terms of the total execution time on CPUs, as shown in Figure 1.
-Therefore, GPUs and accelerators have focused on the optimization of the GEMM-based operators, which significantly enhanced the computational performance (e.g., latency and throughput) of end-to-end ML model inference, as highlighted in Figure 1.
+Therefore, GPUs and accelerators have focused on the optimization of the GEMM-based operators, which significantly enhanced the computational performance (e.g., latency and throughput) of end-to-end ML model inference.
 Figure 1: The latency breakdown into GEMM and non-GEMM operators on AMD EPYC 7763 + NVIDIA A100 GPU.
@@
 Figure 1 shows the profiled latency breakdown into GEMM and non-GEMM operators running inferences on state-of-the-art large language (GPT2-XL [7]) and image classification (Swin Transformer [38]) models.
-The motivational data imply a major shift of the landscape of Amdahl’s law [24] for ML inference acceleration, indicating that we now need to consider non-GEMM operators as one of the major optimization targets in the ML system optimization.
+The motivational data show that the non-GEMM operations now can account for the majority of the latency with GEMM acceleration, indicating that we now need to consider non-GEMM operators as one of the major optimization targets in the ML system optimization.
 However, the research community today lacks a thorough and systematic performance analysis and characterization of non-GEMM operators in the latest models, which hinders the development of non-GEMM oriented optimization techniques.
-Therefore, we collect widely-used ML models from Hugging Face [61] and Torchvision [38] and perform a thorough performance characterization of of non-GEMM operators in the 18 latest models of four major task domains: Image Classification (IC), Image Segmentation (IS), Object Detection (OD), and Natural Language Processing (NLP).
-We evaluate the effect of GPU acceleration on the relative latency across GEMM and non-GEMM operators in collected models and conduct deep-dive analysis on the impact of different hardware platforms (mobile, workstation, and datacenter), deployment software, and common optimizations (operator fusion and quantization).
+Therefore, we collect widely-used ML models from Hugging Face [59] and Torchvision [55] and perform a thorough performance characterization of of non-GEMM operators in the 17 latest models of four major task domains: Image Classification (IC), Image Segmentation (IS), Object Detection (OD), and Natural Language Processing (NLP).
+We evaluate the effect of GPU acceleration on the relative latency across GEMM and non-GEMM operators in collected models and conduct deep-dive analysis on the impact of different hardware platforms (workstation and data center), deployment software, and common optimizations (operator fusion and quantization).
 Based on our case studies, we highlight that the non-GEMM performance challenge is common in accelerated inferences and existing optimization techniques (e.g., operator fusion) cannot completely address the challenges.
@@
 To facilitate such research in non-GEMM-oriented optimization techniques, we build an open-source benchmark specialized in non-GEMM performance anlysis, NonGEMM Bench, which will be released after publication.
-NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch [46], ONNX [7], and TensorRT [43], in addition to the preset of non-GEMM operators collected from the selected 16 popular models, which provides desired flexibility to users for follow-up research.
+NonGEMM Bench can profile arbitrary non-GEMM operators supported by PyTorch [44], ONNX [10], and TensorRT [4], in addition to the preset of non-GEMM operators collected from the selected 17 popular models, which provides desired flexibility to users for follow-up research.
 We summarize our contributions as follows:
@@
   • We analyze the impact of a common non-GEMM-aware optimization, operator fusion and show that operator fusion does not completely mitigate non-GEMM bottleneck for all models, which motivates follow-up research in non-GEMM performance optimizations
-  • We evaluate the performance of non-GEMM operators with quantization and show the non-GEMM bottleneck aggravates with data.
-  • We will open-source NonGEMM Bench, an expandible benchmark flow that enables thorough non-GEMM performance characterization for any model supported by ONNX runtime, Tensor RT, and PyTorch, to facilitate non-GEMM-oriented research.
+  • We evaluate the performance of non-GEMM operators with quantization and quantitatively show the non-GEMM bottleneck aggravates with quantization.
+  • We open-source NonGEMM Bench, an extensible benchmark flow that enables thorough non-GEMM performance characterization for any model supported by ONNX runtime, TensorRT, and PyTorch, to facilitate non-GEMM-oriented research.
 II Background
-Figure 2: A description of Linear (a) and Conv1D (c) operators as GEMM operators example, and that of non-maximum suppression [21] (b) and Layer Normalization [6] (d) as a non-GEMM operator example.
+Figure 2: Descriptions of example non-GEMM and GEMM operators. (a) (Non-GEMM) non-maximum suppression [24], (b) (GEMM) Conv1D (c) (Non-GEMM) Layer Normalization [9], and (d) (GEMM) Linear.
 II-A ML Operators
-ML operators are the building blocks for ML models, which define the computation over one or multiple input tensors.
+ML operators are the building blocks of ML models, which define the computation over one or multiple input tensors.
 Examples include convolution (Conv2d), matrix multiplication (linear, BMM, etc.), activation, and normalization, as listed in Figure 2.
-We categorize operators into two classes: GEMM operation-based and the others, which we term as non-GEMM operators.
+We categorize operators into two classes: GEMM operation-based (”GEMM operators”) and the others (”Non-GEMM operators”).
 We discuss each class of ML operators next.
-II-A1 GEMM-based Operators (GEMM Operators)
+GEMM-based Operators (GEMM Operators).
 GEMM-based operators (or GEMM operators) refer to all the ML operators that can be represented as a matrix multiplication operation, which include linear, Conv2d, and batched-matrix multiplication (BMM).
-For example,  Figure 2 (a) and (c) illustrate two popular GEMM operators: Linear and Conv2d operators, respectively.
+For example,  Figure 2 (b) and (d) illustrate two popular GEMM operators: Linear and Conv2d operators, respectively.
 Each operator can be represented into a perfectly nested loop with multiply-and-accumulate (MAC) operation in the inner-most loop.
 Note that variants that are not matrix multiplication in the default form like Conv2d can be converted into GEMM (e.g., im2col [14]), which motivated the term, GEMM operator.
-GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in in Figure 1 show.
+GEMM-based operators are known to be compute-intensive, which accounts for the majority of the execution unless accelerated by GPUs or accelerators, as CPU results in Figure 1 show.
 However, they have regular computation patterns that can be summarized as a perfectly nested for loop.
 The regular pattern allows various loop optimization techniques such as loop reordering, tiling, and parallelization, which is referred to as dataflow [43, 34, 62] With the dominance of GEMM operators in execution time and high optimization potential together, GEMM operators have been the prime optimization target for acceleration, which led to high-performance GPUs (e.g., H100 [48]) and accelerators [31].
-II-A2 Non-GEMM Operators
+Non-GEMM Operators.
 Non-GEMM operators refer to all ML operators other than GEMM operators.
@@
 Because of their diverse functionalities, their computation patterns are often not a perfectly nested loop with MAC, which can also involve non-linear functions and memory-oriented operations.
-For example,  Figure 2 (b) shows non-maximum suppression (NMS) operator often found in R-CNN model variants [21, 53].
+For example,  Figure 2 (a) shows non-maximum suppression (NMS) operator often found in R-CNN model variants [24, 50].
 As found in the example, the entire operation cannot be summarized into single perfectly-nested loop, which involve other operations such as sort and filtering.
 In addition, the operation involves a conditional statement, which introduces non-deterministic behaviors to the operator.
-The layer normalization example in Figure 2 (b) also shows another key characteristic of the non-GEMM operators: non-linear functions.
+The layer normalization example in Figure 2 (c) also shows another key characteristic of the non-GEMM operators: non-linear functions.
 Because of such characteristics distinguished from GEMM operators, optimization methodologies for GEMM operators cannot be applied to accelerate non-GEMM operators.
-To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 16 recent models in the computer vision and natural language processing domains.
+To understand the extent of the non-GEMM operators, we analyze non-GEMM operators in 17 recent models in the computer vision and natural language processing domains.
 We select models based on their popularity in the Hugging Face to obtain realistic workload.
-We list the models we investigated in Table I.
-Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table II.
+We list the models we investigated in Table II.
+Based on our analysis, we categorize non-GEMM operators based on their functionality and summarize their usage in models and characteristics in Table I.
   • Normalization.
@@
 Activation operators introduce non-linearity into the model.
-Rectified Linear Unit (ReLU) function [41] is an example of activation operators used in CNN based ML models  [54, 22, 55].
-It injects non-linearity into the model based on the sign of the data by applying the function $X_{i}=Max(0,X_{i})$ element-wise to the data, where $X_{i}$ is the i-th data point.
-Another variant of activation operators is the the Gaussian Error Linear Units function (GELU) [23] which is a popular activation function adopted in transformer based ML models s [4, 64, 32, 37].
-Unlike ReLU, GELU accounts for the value of the data when inserting the non-linearity and not only the sign.
-It multiplies the input $X_{i}$ by the Cumulative Distribution Function (CDF) of a Gaussian distribution: $GELU(X_{i})=X_{i}*\phi(X_{i})$ [23] where $X_{i}$ is the i-th data and $\phi$ is the CDF of the Gaussian distribution.
+Rectified Linear Unit (ReLU) function [41] is an example of activation operators widely used in CNN based ML models  [51, 25, 52].
+ReLU injects non-linearity into the model based on the sign of the data by applying the element-wise function, $ReLU(X)=Max(0,X)$.
+Another variant of activation operators is the the Gaussian Error Linear Units function (GELU) [26], which is a popular activation function adopted in transformer based ML models [7, 61, 33, 38].
+Unlike ReLU simply gates out negative values to be 0, GELU requires to compute the Cumulative Distribution Function (CDF) of a Gaussian distribution, which is often denoted as $\phi$.
+GELU multiplies the input $X$ by the Cumulative Distribution Function (CDF) of a Gaussian distribution ($\phi$): $GELU(X)=X*\phi(X)$ [26].
   • Memory Operators.
@@
   • Element-Wise Arithmetic.
-Elemennt-wise arithmetic operators cover non-GEMM operators applied on individual elements in a tensor.
-For example, Figure 3 (c) contains an element-wise division applied to scale the elements of tensors in the attention block
+Element-wise arithmetic operators refer to all the operations applied on individual elements in a tensor (other than activations).
+For example, Figure 3 (c) contains an element-wise division applied to scale the elements of tensors in the attention block.
   • RoI Selection.
-RoI selection operators are found in R-CNN variant detection models.
+RoI selection operators are found in R-CNN variants.
 [50, 24].
@@
 Figure 3: Architectures of three popular ML model families.
-II-B ML Models and Popular Tasks
-The heterogeneity in non-GEMM operators enabled ML developers to build models supporting a wide range of tasks (e.g.
-Computer Vision and NLP).
-As highlighted in  Figure 3, computer vision ( Figure 3 a and b) and NLP ( Figure 3) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
-For example, classical models in image classification shown in  Figure 3 (a) consist of deep convolution networks (CNN).
-which cascade Conv2d (GEMM) with normalization and activation operators [54, 22].
-Object Detection models also rely on convolution network for feature extraction, region proposal and classification.
-However, they pair the CNNs with ROI selection non-GEMM operators to process and filter the detections (Figure 3 (b)).
-On the other hand, NLP and language models rely on the transformer architecture, which leverages the attention mechanism introduced in [59].
-Transformers combine BMMs and Linear GEMM operators with normalization, memory and arithmetic non-GEMM operators as shown in Figure 3.
-This diversity in non-GEMM operators and model architectures across tasks paired the preliminary data presented in  Figure 1 motivate us to study the impact of non-GEMM operators have on the inference performance of popular models.
-Figure 4: An overview of NonGEMM Bench flows
-III Performance Characterization Methodology
-To understand the realistic performance landscape of the latest ML models with non-GEMM workloads, we must (1) capture operator level performances in end-to-end inferences to understand, (2) use widely-used models by the research community and industry, (3) cover diverse task domains, and (4) use real datasets.
-However, ML Benchmarks available today (e.g., MLPerf [52]), unfortunately, do not satisfy all the requirements since they do not focus on the non-GEMM operators.
-Long-tail bench [34] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domain.
-Therefore, to facilitate our non-GEMM operators analysis and better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench.
-NonGEMM Bench provides operator-level breakdown of end-to-end inference latency in the operator graph level, which enables detailed non-GEMM operator performance analysis, as we present in Section IV.
-To capture the performance in the latest ML workload, we select 15 highly downloaded (more than 10K downloads on average) models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench and our analysis.
-We discuss the datasets and models in NonGEMM Bench in detail and describe the structure of NonGEMM Bench next.
-III-A Evaluated Models
-Table I lists the NonGEMM Bench model registry which contains 15 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 0.85M to 7B, demonstrating the diverse model coverage of NonGEMM Bench.
-The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
-III-A1 Computer Vision (CV)
-Deep learning is widely used in the CV domain [32]  [22]  [53]. NonGEMM Bench covers three important tasks in the domain:
-Image Classification (IC).
-Image classification refers to a CV task that identifies a class label of a given image.
-Image classification models extract features (i.e., high-level and dimensional information of the input image) from an input image and report the class label utilizing the features.
-NonGEMM Bench includes eight most popular IC models in HuggingFace [1]: Three variants of Vision Transformer [32], and three variants of Swin Transformer  [37].
-Image Segmentation (IS).
-Image segmentation refers to a computer vision task that identifies the area in a image for each class.
-Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels that belong to each object.
-NonGEMM Bench includes two state-of-the-art IS models: Segformer  [64] and MaskFormer  [10].
-Object Detection (OD).
-Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object.
-OD models extract features and generate region proposals, which refer to the candidate locations and bounding boxes of objects.
-Using Region of Interest (RoI) processing non-GEMM operators in Table II, OD models refine raw region proposals generated by a region proposal network.
-The refined RoIs are used as inputs to the CNN classifier at the end, and the classifier determines the class label of objects inside each refined RoI.
-NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [53], MaskRCNN  [21], and DETR  [9].
-III-A2 Natural Language Processing (NLP)
-NLP refers to tasks involving the analysis and understanding of human (natural) language.
-NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [68] [8].
-Transformer [59] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state-of-the-art generative large language models like GPT  [50] and Llama  [58].
-Figure 3 (c) shows the layer architecture of GPT’s transformer.
-It consists of a self-attention block built by cascading GEMM operators with Normalization, Memory, Logit Comuputation and Element-wise Arithmetic non-GEMM oeprators ( Table II).
-NonGEMM Bench includes five popular NLP models  [3]: Bert [31], three variants of GPT2 [4], and Llama2-7b [58].
-III-B Profiling Methodology
-To enable the thorough analysis of the landscape of non-GEMM performance presented in Section IV, NonGEMM Bench provides an end-to-end profiling flow to breakdown the inference latency at the operator level.
-We discuss input, output, and the benchmark software structure.
-III-B1 NonGEMM Bench Inputs
-NonGEMM Bench receives target deployment software flow (PyTorch, ONNX runtime, or TensorRT), dataset, and other configurations listed in Figure 4 as inputs.
-Optionally, users can provide custom models with associated datasets to profile their workloads.
-By default, NonGEMM Bench runs the 15 models in Table I.
-We describe each input in detail as follows.
-Models.
-As described in  Section III-A, NonGEMM Bench includes a registry of 15 selected popular ML models.
-Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators.
-Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
-Deployment Flow. NonGEMM Bench deploys its models using three popular inference frameworks: ONNX Runtime [17], PyTorch [46] TensorRT [43].
-Datasets.
-To evaluate the models, NonGEMM Bench utilizes real data from popular datasets in each domain.
-We use ImageNet 2012  [15] and MS COCO  [36] for computer vision tasks.
-As for language models, we use wikitext dataset [39] available on HuggingFace.
-For custom models, users can specify their own dataset for the model.
-Misc. Configurations. This includes the batch size, the number of profiling iterations, and the target device.
-TABLE I: Tasks and Models Evaluated in Section IV
-Application Models # Params Dataset
-Image Classification (IC) ViT base (Vt-b) [32] 307M ImageNet  [15]
-ViT large (Vt-l) [32] 307M
-ViT huge (Vt-h) [32] 632M
-Swin tiny (Sw-t) [37] 29M
-Swin small (Sw-s) [37] 50M
-Swin base (Sw-b) [37] 88M
-Object Detection (OD) FasterRCNN (FRCNN) [53] 42M COCO  [36]
-MaskRCNN (MRCNN) [21] 44M
-DETR [9] 41M
-Image Segmentation (IS) Maskformer (MF) [10] 102M COCO  [36]
-SegFormer (Seg) [64] 3.7M
-Natural Language Processing (NLP) GPT2 [4] 117M wikitext  [39]
-GPT2 Large (gpt2-l) [4] 762M
-GPT2 X-Large (gpt2-xl) [4] 1.5B
-Llama 2-7B [58] 7B
-Bert [31] 110M
-III-B2 End-to-End Inference Flow
-The end-to-end inference flow profiles NonGEMM Bench models running inferences on data samples from the selected dataset.
-First, the Graph Extractor extracts the operator graph of a given model based on the selected deployment flow.
-NonGEMM Bench calls graph exporters from the HuggingFace transformers library [62] or from the PyTorch libraries.
-Second, model specific Data Preprocessing functions fetch the raw data from the specified dataset, clean the data, and apply tokenization or image to tensor transformations to generate input tensors for the profiling stage.
-Third, End-to-End Inference Profiling functions run the profiling iterations on the received the model and its inputs.
-The profiling functions are deployment flow specific.
-For ONNX Runtime, we invoke Execution Provider profiling in the inference session.
-Currently, we only support operator level profiling for CPU and CUDA execution providers [17] because profiling support to other ONNX RUNTIME backends is not stable yet.
-For PyTorch, we use the PyTorch Profiler [49].
-For TensorRT, we leverage the TensorRT open source software framework profiling APIs [44].
-III-B3 NonGEMM Bench Outputs
-NonGEMM Bench generates many statistics organized into three categories: performance, workload, and non-GEMM-specific reports.
-The performance report includes key performance metrics such as the end-to-end latency with operator level break-downs and peak memory usage.
-The workload report includes the types of operators and the shape of the tensors for each operator captured during inference on realistic data.
-Finally, the non-GEMM report provides insights on non-GEMM operators, such as the number of operator variants of the same class of non-GEMM operator (e.g., layernorm implementation could have multiple variants for each ML framework) and non-GEMM operator trace on different domains.
-TABLE II: Non-GEMM operators in eight selected model variants from Table I and their characteristics.
+TABLE I: Non-GEMM operators in eight selected model variants from Table II and their characteristics. Example input shapes are captured based on inferences using real datasets.
 Operator Group Operator Model Single Operation Single Operand Non Linearity Dynamicity Reduction Example Input Shape
@@
 Interpolation Interpolate Segformer [2, 256, 128, 128]
+II-B ML Models and Popular Tasks
+The heterogeneity in non-GEMM operators enabled ML developers to build models supporting a wide range of modalities and tasks (e.g.
+computer vision and NLP).
+As highlighted in  Figure 3, computer vision (Figure 3 (a) and (b)) and NLP (Figure 3 (c)) models are characterized by distinct architectures leveraging unique combinations of GEMM and non-GEMM operators.
+For example, traditional image classification models are often based on the convolutional neural network (CNN) architecture, which cascades GEMM (Conv2d) and non-GEMM (normalization and activation)operators [51, 25].
+Object detection models, such as Mask R-CNN [24], often utilize CNNs for feature extraction, region proposal, and classification, as illustrated in Figure 3 (b).
+Unlike image classification models, they combine the CNNs with unique non-GEMM operators such as non-maximum suppression (NMS) and ROI Align to process and filter the bounding boxes for objects.
+On the other hand, recent language models employ the transformer architecture, which leverages the attention mechanism introduced in [57].
+Transformers combine a unique set of GEMM (BMMs and Linear) and non-GEMM (normalization, memory, and element-wise arithmetic) operators, as shown in Figure 3 (c).
+As we can find in the aforementioned examples, the model architectures and the combination of GEMM and non-GEMM operators are diverse.
+This would mean that the performance implication of non-GEMM operators would vary across models, as our motivational data presented in  Figure 1 show.
+This motivates a thorough characterization study that investigates (1) if the non-GEMM performance challenge is pervasive across popular models and (2) how significant their implication is, under widely adopted optimization techniques (e.g., integer quantization [29] and operator fusion [42]).
+Therefore, we conduct a thorough case study of the non-GEMM performance horizon.
+Figure 4: An overview of NonGEMM Bench flow.
+III Performance Characterization Methodology
+To understand the realistic performance landscape of the latest ML models with non-GEMM workloads, we must (1) capture operator level performances in end-to-end inferences, (2) use widely-used models by the research community and industry, (3) cover diverse task domains, and (4) use real datasets.
+However, ML Benchmarks available today (e.g., MLPerf [47]), unfortunately, do not satisfy all the requirements since they do not focus on the non-GEMM operators.
+Long-tail bench [35] identified a similar problem as this work, but it focuses on a limited set of custom kernels, which fails to represent broad task domains.
+Therefore, to facilitate our non-GEMM operators analysis and better-understand the impact of non-GEMM operators on system performance, we develop a new ML benchmark, NonGEMM Bench.
+NonGEMM Bench provides operator-level breakdown of end-to-end inference latency in the operator graph level, which enables detailed non-GEMM operator performance analysis, as we present in Section IV.
+To capture the performance in the latest ML workload, we select 17 highly downloaded (more than 10K downloads on average) models from HuggingFace [1] to enhance the representativeness of NonGEMM Bench and our analysis.
+We discuss the models and datasets adopted in NonGEMM Bench in detail and describe the structure of NonGEMM Bench next.
+III-A Models included in NonGEMM Bench
+Table II lists the NonGEMM Bench model registry which contains 17 models based on state-of-the-art CNN and Transformer architectures with number of parameters ranging from 3.7M to 7B, demonstrating the diverse model coverage of NonGEMM Bench.
+The selected models cover four major task domains in ML, which include Image Classification (IC), Object Detection (OD), Image Segmentation (IS), and Natural Language Processing (NLP).
+Image Classification (IC).
+Image classification refers to a CV task that identifies a class label of a given image.
+Image classification models extract features (i.e., high-level and dimensional information of the input image) from an input image and report the class label utilizing the features.
+NonGEMM Bench includes six most popular IC models in HuggingFace [1]: Three variants of Vision Transformer [33], and three variants of Swin Transformer  [38].
+Image Segmentation (IS).
+Image segmentation refers to a computer vision task that identifies the area in a image for each class.
+Like IC models, IS models also extract features and utilize them to identify objects located in an image and spatially separate them by highlighting pixels that belong to each object.
+NonGEMM Bench includes two state-of-the-art IS models: Segformer  [61] and MaskFormer  [13].
+Object Detection (OD).
+Object detection refers to a computer vision task that identifies the location of objects in an image and outputs the bounding box of each object.
+OD models extract features and generate region proposals, which refer to the candidate locations and bounding boxes of objects.
+Using Region of Interest (RoI) processing non-GEMM operators in Table I, OD models refine raw region proposals generated by a region proposal network.
+The refined RoIs are used as inputs to the CNN classifier at the end, and the classifier determines the class label of objects inside each refined RoI.
+NonGEMM Bench includes three popular OD models [2]: FasterRCNN  [50], MaskRCNN  [24], and DETR  [12].
+Natural Language Processing (NLP).
+NLP refers to tasks involving the analysis and understanding of human (natural) language.
+NLP models extract context and features from an input text sequence and use the extracted context and features to perform multiple applications like translation and text generation [64] [11].
+Transformer [57] based DNNs have become the dominant model architecture in NLP and are the backbone of popular state-of-the-art generative large language models like GPT  [46] and Llama  [56].
+Figure 3 (c) shows the layer architecture of GPT’s transformer.
+It consists of a self-attention block built by cascading GEMM operators with Normalization, Memory, Logit Computation and Element-wise Arithmetic non-GEMM operators ( Table I).
+NonGEMM Bench includes six popular NLP models [3]: Bert [19], three variants of GPT2 [7], Llama2-7b [56], and Mixtral 8x7B [30].
+III-B NonGEMM Bench Inputs
+As shown in Figure 4, NonGEMM Bench flow receives workload and dataset information (default: 17 NonGEMM Bench models in Table II), target deployment flow, and other configurations such as the batch size and number of runs for performance characterization.
+Models.
+As described in  Section III-A, NonGEMM Bench includes a registry of 17 selected popular ML models.
+Nonetheless, we designed NonGEMM Bench to be easily expandable to accommodate rapidly evolving ML models that constantly introduce new operators.
+Users can benefit from the features of our benchmark by simply plugging their new models into the NonGEMM Bench model registry (Figure 4) by specifying the model class, its weights and any data preprocessor.
+Deployment Flow. NonGEMM Bench supports four popular inference frameworks: ONNX Runtime [49], PyTorch [44], TensorRT [4], and TorchInductor [8].
+Datasets.
+To evaluate the models, NonGEMM Bench utilizes real datasets popular in each domain.
+We use ImageNet 2012  [17] and MS COCO  [37] for computer vision tasks.
+As for language models, we use wikitext dataset [39] available on HuggingFace.
+For custom models, NonGEMM Bench allows users to specify their own dataset for their models.
+Configurations.
+Users can specify detailed configurations for the performance characterization using NonGEMM Bench.
+The configurations include the batch size, the number of profiling iterations, and the target device.
+III-C NonGEMM Bench Outputs
+NonGEMM Bench generates many statistics organized into three categories: performance, workload, and non-GEMM-specific reports.
+Performance Report. The performance report includes key performance metrics such as the end-to-end latency with operator level break-downs (Figure 6) and the end-to-end energy consumption (Figure 5).
+Workload Report. The workload report includes the types of operators and the shape of the tensors for each operator captured during inference on realistic data.
+Non-GEMM Report.
+The non-GEMM report provides insights on non-GEMM operators, such as the number of operator variants of the same class of non-GEMM operator (e.g., DETR [12] employs two variant of BatchNorm, a custom implementation and one in the PyTorch operator library) and non-GEMM operator trace on different domains.
+TABLE II: Tasks and Models Evaluated in Section IV
+Application Models # Params Dataset
+Image Classification (IC) ViT base (Vt-b) [33] 307M ImageNet  [17]
+ViT large (Vt-l) [33] 307M
+ViT huge (Vt-h) [33] 632M
+Swin tiny (Sw-t) [38] 29M
+Swin small (Sw-s) [38] 50M
+Swin base (Sw-b) [38] 88M
+Object Detection (OD) FasterRCNN (FRCNN) [50] 42M COCO  [37]
+MaskRCNN (MRCNN) [24] 44M
+DETR [12] 41M
+Image Segmentation (IS) Maskformer (MF) [13] 102M COCO  [37]
+SegFormer (Seg) [61] 3.7M
+Natural Language Processing (NLP) GPT2 [7] 117M wikitext  [39]
+GPT2 Large (gpt2-l) [7] 762M
+GPT2 X-Large (gpt2-xl) [7] 1.5B
+Llama 2-7B [56] 7B
+Bert [19] 110M
+Mixtral 8x7B [30] 46.7B
+III-D NonGEMM Bench Performance Characterization Flow
+NonGEMM Bench’s software flow accepts inputs described in Section III-B and generates outputs described in Section III-C.
+Internally, NonGEMM Bench includes graph extractor, data preprocessing, and performance profiling modules.
+Graph Extractor.
+The Graph Extractor module extracts the operator graphs of input models based on the selected deployment flow.
+NonGEMM Bench utilizes graph exporters in the HuggingFace Transformers [59] and PyTorch.
+Data Preprocessing.
+The Data Preprocessing module includes model-specific preprocessing functions that fetch raw data from the target dataset, clean the data, and apply desired transformations (e.g., tokenization and image to tensor).
+Performance Profiling.
+The Performance Profiling (PP) module launches the inferences, collects performance statistics, and generates output reports discussed in Section III-C.
+The module selects appropriate profiling functions based on the deployment flow choice.
+For PyTorch, the PP module utilizes the PyTorch Profiler [5].
+For TensorRT [6], the PP module leverages its profiling APIs.
+For ONNX Runtime [49], the PP module invokes Execution Provider profiling.
 IV Case Studies
-We perform a thorough performance analysis of models in Table I on multiple hardware platforms listed in  Table III.
-The evaluated hardware platforms span workstation and data center-class CPUs and GPUs.
-We employ PyTorch, ONNX runtime, and TensorRT as deployment flows to investigate their impact on the performance horizon.
-Using the aforementioned workload, hardware platforms, and deployment flow, we characterize the GEMM/non-GEMM performance for each model in NonGEMM Bench and investigate the impact of the hardware platform and deployment flow.
-In addition, we also investigate the impact of widely adopted optimizations (operator fusion and quantization) and show that the non-GEMM performance is an issue even after applying common optimizations.
+We conduct a thorough performance analysis of models in Table II on a data center and a workstation-class platform, as listed in  Table III.
+We employ PyTorch for our the main performance characterization and use ONNX Runtime, TensorRT, and TorchInductor for deep-dive studies (e.g., the impact of deployment flow choice and operator fusion).
+We first focuse on the GEMM and non-GEMM performance horizon (Section IV-A).
+Also, we provide deeper insights into the non-GEMM performance horizon by investigating the impact of deployment flow and operator fusion (Section IV-B), and the impact of quantization (Section IV-C).
 TABLE III: Hardware platform configurations used for case studies.
@@
 IV-A Non-GEMM Performance Characterization Results
-We first conduct a performance characterization study using PyTorch and present the characterization results on datacenter and workstation platforms (Platforms A and B in Table III) in Figure 5.
-We focus on those two platforms since some models (e.g, large language models) do not fit into the mobile devices.
-As expected, the relative contribution of non-GEMM operators to the end-to-end latency significantly increases after GPU acceleration, on average, from 17.2% to 42.3%.
+Figure 5: End-to-End inference GPU energy consumption of models running on the Data Center (CPU + GPU) configuration.
+We conduct a performance characterization study using PyTorch and present the results in Figure 6.
+Aligned with what we observed in  Figure 1, the relative contribution of non-GEMM operators to the end-to-end latency significantly increases after GEMM acceleration using a GPU, from 17.2% to 42.3%, on average.
 However, we observe each model show different trends, mainly affected by the non-GEMM operator types and the number of GEMM and non-GEMM operators in the original model.
@@
 MaskFormer utilizes Swin Transformer as its backbone, which introduces many memory operators as we discussed for Swin Transformer.
-As a result, memory operator is the dominant non-GEMM operator, which accounts for 40.8% of the total latency, on average, as we can observe in Figure 5 (f) and (n).
+As a result, memory operator becomes the most dominant non-GEMM operator, which accounts for 40.8% of the total latency, on average, as we can observe in Figure 6 (f).
 NLP Task: GPT2.
-The latencies of non-GEMM operators in GPT2 variants are considerable, which account for 45.0%, on average, as we observe in Figure 5 (h) and (p).
-The dominant non-GEMM operator is activation, GELU, which required 26.4% of the total latency.
+As we observe in Figure 6 (h) for both platforms, the latency of non-GEMM operators in GPT2 variants is considerable, which account for 45.0%, on average.
+The most dominant non-GEMM operator is an activation function, GELU, which accounts for 26.4% of the total latency.
+Summary.
+We observe GPU acceleration significantly increases the percentage of non-GEMM operators in the end-to-end latency, which amplifies the importance of non-GEMM operators in the performance optimization process.
+Also, we observe the most dominant non-GEMM operators are diverse depending on the model.
+The results indicate that an optimization technique tailored for a single operator cannot fully address the non-GEMM performance challenge, which motivates a holistic optimization approach for wide non-GEMM operators or a balanced specialization for a set of non-GEMM operators in a target workload.
 TABLE IV: Most time-consuming non-GEMM operator groups for selected Models (Platform A, with GPU acceleration, average across batch sizes).
@@
 bert Norm 13.1
-Summary.
-We observe GPU acceleration significantly increase the percentage of non-GEMM operators and increase the importance of non-GEMM operators in performance optimization.
-Also, the dominant operators are diverse in each model, which indicates an optimization technique tailored for a single operator cannot fully address the non-GEMM performance problem.
-Figure 5: Breaking down the execution time of models running on the Data Center (CPU only) and (CPU + GPU) configurations across operator groups.
+Mixtral Memory 43.1
+Figure 6: Latency break-downs of NonGEMM Bench models into the operator granularity. We show CPU-only (left column) and CPU+GPU (right column) results on two evaluated platforms listed in Table III.
 IV-B The Impact of Deployment Flow on Non-GEMM Performance
 Deployment flows such as ONNX Runtime [49] and TensorRT [4] are widely used for serving model inferences.
-Such flows apply various optimizations to each model, which includes the computational graph level optimizations (e.g., operator fusion) and backend assignment (e.g., utilizing Tensor Core in Nvidia GPUs).
-To understand the effect of deployment frameworks on the operator performance, we conduct two case studies (1) comparing PyTorch and ONNX Runtime (ORT) results on LLMs(focus: general optimizations w/o operator fusion) on LLMs, and (2) comparing PyTorch and TensorRT results (focus: operator fusion).
-IV-B1 Case Study 1: Non-GEMM Performance on LLMs across PyTorch and ORT
-We run two LLMs (GPT2 and Llama2) using the CUDA execution provider in ORT and compare their non-GEMM performance percentage in Figure 6.
-In the results, we observe the presence of non-GEMM operators significantly increase, from 52.6% to a 80.75%, on average.
-We observe the percentage of memory operators significantly amplifies after applying ORT, from 3.2% to 66.8%.
-Such a result originates from the limitation in operator support in ORT.
-Many memory operators in the evaluated LLMs are not supported by the cuda execution provider in ORT, which leads to inefficient execution on CPUs involving data transfer between a CPU and a GPU.
-Combined with the high frequency of those operators, the inference latency of memory operators increases significantly, as shown in Figure 6 (b).
+Such flows apply various optimizations to each model, which includes the computational graph optimizations (e.g., operator fusion) and backend assignment (e.g., utilizing Tensor Core in Nvidia GPUs).
+To understand the impact of deployment frameworks on the non-GEMM operator performance, we conduct two case studies: (1) comparing PyTorch and ONNX Runtime (ORT) results on LLMs (focus: general optimizations w/o operator fusion) and (2) comparing PyTorch, TorchInductor, and TensorRT results (focus: operator fusion).
+[Case Study 1] Non-GEMM Performance on LLMs across PyTorch and ORT.
+We profile the GEMM and non-GEMM performance of two LLMs (GPT2 and Llama2) on the platform A, using the CUDA execution provider in ORT.
+As the results presented in Figure 7 show, we observe the presence of non-GEMM operators significantly increase, from 52.6% to a 80.75%, on average.
+We observe the percentage of memory operators significantly increases if we switch from PyTorch to ORT, from 3.2% to 66.8%, although the absolute end-to-end latency decreases.
+Such a result originates from ORT’s significant performance boost of other operators (Lllma2) and ORT’s limited efficiency on memory operators (GPT2-XL).
+Many memory operators in the evaluated LLMs are not supported by the CUDA execution provider in ORT, which leads to inefficient execution on CPUs involving costly data transfer between a CPU and a GPU.
+Combined with the high frequency of such operators, the relative contribution of memory operators to the end-to-end inference increases significantly, as shown in Figure 7 (b).
 The results imply two insights: (1) Model deployment flow can significantly aggravate the non-GEMM performance challenge and (2) the dominant non-GEMM operators differ depending on the operator support in a deployment flow.
 Figure 7: The impact of deployment software toolchain into the latency breakdown on language models. (a) PyTorch and (2) ONNX Runtime on a data center class GPU (A100).
-Figure 7: The latency of non-GEMM operator in (a) Swin-b, (b) DETR, and (c) Segformer is still considerable despite applying operator fusion in TensorRT.
-The numbers at the top of each bar represent the latencies in milliseconds.
-IV-B2 Case Study 2: Non-GEMM Performance with Operator Fusion
-Operator fusion is a key optimization technique for accelerating inference workloads [14, 25, 57, 42, 43, 30].
-Fusion combines multiple operators in a single kernel to reduce the number of costly kernel launches and minimize the number of redundant offchip memory accesses [42].
+Figure 8: The latency of non-GEMM operator in (a) Swin-t, (b) Swin-b, (c) DETR, and (d) Segformer is still considerable despite applying operator fusion in TorchInductor (middle) and TensorRT (right).
+[Case Study 2] Non-GEMM Performance with Operator Fusion.
+Operator fusion is one of the key optimization technique for accelerating inference workloads [16, 27, 54, 42, 4, 32].
+Operator fusion combines multiple operators in a single kernel to reduce the number of costly kernel launches and minimize the number of redundant offchip memory accesses [42].
 TensorRT [4] is a widely adopted inference framework released by Nvidia that applies the operator fusion technique targeting GPUs.
 Operator fusion in TensorRT detects specific patterns (e.g., three consecutive element-wise operators [4]) in the operator graph and fuses nodes captured in the patterns to enhance inference performance by reducing redundant memory accesses around non-GEMM operators.
-Impact of Operator Fusion.
-To understand the impact of operator fusion on non-GEMM performance horizon, we conduct a case study on four models, which compares the TensorRT (with fusion) and PyTorch (without fusion).
-We present the results in Figure 7, which shows the inference latency breakdown between GEMM and non-GEMM operators.
-We notice that fusion does not completely address the non-GEMM bottleneck in Swin-b model.
-For example, the percentage of non-GEMM operators in the total latency changes from 56.2% to 32.2%, on average, after operator fusion by TensorRT.
-The reduction in the percentage is based on the improvement of non-GEMM performance via operator fusion, which reduces 88.6% of latency after fusion, as summarized in Table V.
-However, the non-GEMM operators still account for 32.2% of total latency, which is considerable.
-This shows that operator fusion cannot eliminate the non-GEMM performance challenge and motivates further studies toward non-GEMM performance.
-Factors Affecting the Effectiveness of Fusion.
-Although most results indicate considerable impact of non-GEMM even after operator fusion, we observe operator fusion on DETR model is exceptionally effective.
+To understand the impact of operator fusion on the non-GEMM performance, we conduct a case study on four models listed in Figure 8, comparing TensorRT (with fusion) and PyTorch (without fusion).
+We present the results in Figure 8, which shows the inference latency breakdown between GEMM and non-GEMM operators on Platform A (data center class).
+The results indicate that fusion mitigates the non-GEMM bottleneck, but it does not completely address the challenge.
+For example on Swin-b, after applying operator fusion by switching to TensorRT from PyTorch, the contribution of non-GEMM operators to the total latency changes from 56.4% to 32.2%, on average.
+The reduction in the percentage is based on the non-GEMM performance improvement via operator fusion, which reduces 88.6% of latency, as summarized in Table V.
+However, the non-GEMM operators still account for 32.2% of total latency.
+This shows that operator fusion cannot eliminate the non-GEMM performance challenge and motivates further studies toward non-GEMM performance optimization.
+Although most results indicate a considerable impact of non-GEMM even after operator fusion, we observe that TensorRT operator fusion on the DETR model is exceptionally effective.
 Therefore, we conduct a deep-dive study, investigating the percentage of fused non-GEMM operators (i.e., fusion rate) and performance improvements after fusion, as listed in Table V.
 We observe the strong non-GEMM performance improvements for DETR originates from high fusion rate of 30%, which led to 13.5$\times$ non-GEMM speedup.
-This leads to the large percentage reduction of non-GEMM in the total latency, from 67% to 18.5%, on average.
+This leads to the large percentage reduction of non-GEMM in the total latency, from 66.5% to 18.5%, on average.
 However, the fusion rate is not the only factor that determines the non-GEMM speedup.
-For example, DETR and Segformer have similar fusion rates (30% andn 27%, respectively), but the amount of non-GEMM performance improvements are significantly different: 13.5$\times$ and 2.39$\times$, respectively.
-We analyze the execution trace and find that batch normalization is the main source of the differences.
-The main difference is the fusion pattern: most batch normalization operators (100% of total) in DETR were fused with GEMM-operators (CONV+BN+ReLu pattern) while those in Segformer were fused with other non-GEMM operators (97.8% of total).
+For example, DETR and Segformer have similar fusion rates (30% and 27%, respectively), but the amount of non-GEMM performance improvements are significantly different: 13.5$\times$ and 2.39$\times$, respectively.
+We analyze the execution trace and identify the fusion pattern around batch normalizations as the main source of the difference.
+Most batch normalization operators (100% of total) in DETR were fused with GEMM-operators (CONV+BN+ReLu pattern) while those in Segformer were fused with other non-GEMM operators (97.8% of total).
 The results indicate that the effectiveness of operator fusion relies on not only the overall fusion rate but also the fusion patterns.
 Our observation confirms that the operator fusion cannot fully address the non-GEMM performance challenge, even if it can be very effective on some patterns.
-IV-C Non-GEMM Performance with Quantization
+TABLE V: The non-GEMM latency before and after applying fusion with TensorRT. The values between brackets represent the percentage with resect to the total inference latency.
+Model Non-GEMM Non-GEMM Latency
+Fusion Rate Before Fusion After Fusion
+Swin-t 8.8% 7.53 ms (56.4%) 0.97 ms (39.0%)
+Swin-b 7.0% 14.59 ms (56.4%) 1.65 ms (32.3%)
+DETR 30.0% 32.17 ms (66.4%) 2.38 ms (18.5%)
+Segformer 27.0% 5.57 ms (41.0%) 2.33 ms (41.0%)
+IV-C The Impact of Quantization Non-GEMM Performance
 Quantization refers to the model optimization technique, which reduces the bit precision of model weights and/or activations to enhance the computational performance and efficiency of DNN inference.
 Quantization is a widely-adopted technique [29, 40] including heavy models like LLMs [21, 36, 60, 18].
-For example, LLM.int8() [16] introduced an 8-bit - 16-bit mixed precision quantization scheme for LLMs.
 LLM.int8() is a state-of-the-art quantization method, which quantizes more than 99% of the linear layers in OPT LLM to an 8-bit precision.
-Therefore, we adopt LLM.int8() and characterize GEMM and non-GEMM performance to understand the impact of quantization on the non-GEMM operator performance problem.
-We conduct a case study on quantization using LLM.int8() on Llama3 LLM on the datacenter platform with GPU acceleration and present the results in Figure 8.
-In the results, we observe that non-GEMM operators dominate the latency after quantization, changing from 29.3% to 76.7%, on average.
+Therefore, we adopt LLM.int8() and characterize GEMM and non-GEMM performance of Llama3 on Platform A to understand the impact of quantization on the non-GEMM operator performance problem.
+As the results in Figure 9 show, we observe that non-GEMM operators dominate the latency after quantization, changing from 29.3% to 76.7%, on average.
 Such a significant shift in the latency distribution is mainly based on the GEMM performance improvements from 8-bit arithmetic, which reduced the latency by 38.2%, on average.
 However, based on our analysis on the execution traces, non-GEMM performance aggravates because the 8-bit data need to be dequantized and re-quantized for non-GEMM operators, which requires 16-bit floating point arithmetic.
-This led to 6510 additional non-GEMM operators into the computational graph, which led to a significant increase of the non-GEMM latency by 5.6$\times$ after quantization.
+This introduced 6510 additional non-GEMM operators into the computation graph, which led to a significant increase of the non-GEMM latency by 5.6$\times$ after quantization.
 Combined together, the overall percentage of non-GEMM operators in the total latency dominate after quantization, which makes non-GEMM operators as the major optimization target.
@@
 As current trends in the LLMs are toward longer sequence lengths [20, 7], the non-GEMM performance issues in longer sequences originating from DQRQ costs will aggravate, which motivates efforts in non-GEMM performance optimization.
-Figure 8: Non-GEMM operators latency contribution to the inference of a quantized Llama3 8B using 8 bit quantization [16]. The numbers at the top of each bar represent the latencies in milliseconds.
-TABLE V: The non-GEMM latency before and after applying fusion with TensorRT. The values between brackets represent the percentage w.r.t the total inference latency.
-Model Non-GEMM Non-GEMM Latency
-Fusion Rate Before Fusion After Fusion
-Swin-t 8.75% 7.53 ms (56.4%) 0.97 ms (39%)
-Swin-b 7% 14.59 ms (58%) 1.65 ms (32.25%)
-DETR 30% 32.17 ms (66.4%) 2.38 ms (18.5%)
-Segformer 27% 5.57 ms (41%) 2.33 ms (41%)
+Figure 9: GEMM and Non-GEMM latency breakdown of the inference latency on Platform A, running an 8-bit quantized Llama3 8B [18].
 IV-D Key Observations and Insights
@@
   • Operator fusion cannnot fully address non-GEMM performance challenge: Although it can significantly improve non-GEMM performance, but its effectiveness heavily depends on the model.
-  • Operator support in deployment flows significantly affects the non-GEMM performance horizon.
+  • Operator support in deployment flows significantly affects the non-GEMM performance.
   • Quantization significantly aggravates the non-GEMM performance challenge due to the imbalanced speedup across GEMM and non-GEMM and quantization/dequantization costs around non-GEMM operators
 V Related Works
-Previous works characterized ML workloads and non-GEMM operators.
-However, the main limitations are listed below: State-of-the-art inference benchmarks  [20, 52] do not thoroughly analyze the performance of Non-GEMM operators.
-Other works characterizing non-GEMM operators focus on a specific application domain [48, 47, 65, 34] for their non-GEMM study.
-In [48],  [27], and  [47] the authors analyze the performance of non-GEMM operators, however they focus on a specific class of transformer language models.
-Similarly, [34] only studies a limited number of operators from computer vision models in a microbenchmarking flow.
-Previous works that also studied non-GEMM operator performance bottlenecks [12, 56, 60, 51], mostly focused on softmax operators in transformers for language models.
-Our work extends on the pioneering efforts of these works by studying non-GEMM operators from 15 models covering 4 various task domains.
+ML Inference Benchmarks.
+Many end-to-end inference benchmarks [23, 47] do not capture operator level performance breakdowns.
+MLPerf Inference [47], an industry standard inference benchmark, offers a flexible and standardized framework to evaluate the performance of inference systems.
+MLPerf Inference framework defines performance metrics and workloads, and supports measuring the performance of realistic inference scenarios across a wide range of software and hardware systems.
+Nevertheless, MLPerf does not offer any operator-level fine-grained latency breakdowns, which makes it unsuitable for understanding the implications of non-GEMM operators on the inference performance.
+Our work provides fine-grained operator level latency breakdowns to understand the impact of non-GEMM operators on the end-to-end inference performance.
+Non-GEMM Characterization.
+Previous works [35, 45, 15, 53, 58, 22] investigate non-GEMM operators, however their characterization and optimization focus on a limited set of operators and applications.
+Longtail bench [35] proposes a microbenchmark specific to a limited set of non-GEMM operators from selected computer vision models.
+It profiles operators without a compute library implementation in a standalone setting, using randomly generated data.
+Because Longtail bench is a microbenchmark suite, it cannot be used to capture realistic interplay between GEMM and non-GEMM operators in real models, which provides insights for inter-operator optimizations.
+In addition, Longtail bench does not provide general insights on non-GEMM performance because it focuses on a specific computer vision application.
+Our work extends on these efforts by studying the non-GEMM performance of 17 popular models in realistic end-to-end inference scenarios covering various task domains.
+Tandem Processor [22] highlights the importance of non-GEMM operator-oriented optimization in ML inference, and proposes a co-processor architecture to mitigate the non-GEMM overhead.
+Tandem Processor characterized the non-GEMM performance in 7 models and identify non-GEMM operators as the emerging bottleneck after accelerating GEMMs.
+Although Tandem is a pioneering work in non-GEMM optimization, our work provides additional and broader insights beyond it.
+Our work evaluates 17 widely adopted models across task domains and offers detailed case studies analyzing the impact of widely adopted optimization techniques, operator fusion and quantization, on the non-GEMM performance.
 VI Conclusion
-Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law to have more presence of non-GEMM operators in the performance horizon.
-In this paper, we conducted a thorough performance analysis of non-GEMM operators in the latest models in various task domains and platforms.
-The results confirm the increasing importance of non-GEMM performance and show model optimizations we apply (e.g., quantization) can significantly aggravate the non-GEMM performance challenge.
+Accelerating GEMM operators in ML inference have changed the major bottleneck from GEMM to non-GEMM operators.
+To understand the latest GEMM/non-GEMM performance landscape with GEMM acceleration, we conducted a thorough performance analysis of non-GEMM operators in the latest models in various task domains and platforms.
+The results confirm the increasing importance of non-GEMM performance and show that common model optimizations (e.g., quantization) can significantly aggravate the non-GEMM performance challenge.
 The dominant non-GEMM operators are diverse across models, which indicates that a specialized optimization targeting a specific operator cannot solve the non-GEMM performance challenge.
@@
 Our performance anlysis results imply that now we need to consider non-GEMM operators as a major optimization target and develop new hardware and software techniques to optimize non-GEMM performance.
-The non-GEMM profiling software we used in this study, NonGEMM Bench, will be released as open-source software, which will contribute to such follow-up research.
+The non-GEMM profiling software we used in this study, NonGEMM Bench, will be released as open-source software, which will contribute to the follow-up research for non-GEMM optimization.